### v2

# Quantum mechanics provides the physical basis of teleological evolutions

Abstract
We show that the quantum computational speedup is due to the teleological character of quantum algorithms, their being evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again). We also show that, under the quantum cosmological assumption and for the Fine-tuned Universe version of the Anthropic Principle, the physical basis of the teleological character of quantum algorithms applies as well to the evolutions of the living for which the teleological notion was originally conceived.
1 Introduction
Let us briefly examine the very long history of the teleological notion—attraction by a final cause. Plato and Aristotle used it to explain the functional structure of the living (Aristotle: teeth were created for the purpose of chewing); of course, at the time, there was no notion of natural evolution. With the advent of this notion, it was applied to the evolutions of the living by a school of natural philosophy. It was eventually dismissed by science for the alleged absence of a physical basis. All this was before the advent of quantum mechanics. Today, the situation of the teleological notion is somewhat complex. To the best of our knowledge, outside the present work, there are no physical researches on it. Let us round things up by saying that in both physics and the natural sciences there is a de facto disbelief of the retrocasusality and finalism implied in the notion.
In this work, we shall show that the relatively recent discovery of the quantum computational speedup should change this de facto position of science. This is because, in the first place, quantum algorithms turn out to be teleological evolutions: indeed, evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again). Since showing this will be surprisingly simple, the fact that, forty years after the discovery of the quantum speedup, it has never been noticed before could be ascribed to that de facto position of science.
We, with coauthors, already went against this position in the evolutionary, retrocausal, approach to the quantum computational speedup [1, 2, 3, 4, 5, 6]. In it, see [4], we showed that all the quantum algorithms examined, comprising the major ones, obeyed the following rule:
In quantum algorithms, it is as if the problem-solver always knew in advance, before beginning her problem-solving action, any of the possible halves of the information that specifies the solution of the problem she will produce and measure in the future and could use this knowledge to produce the solution with fewer computation steps. The number of computation steps needed in the optimal quantum case is that needed in the optimal logical case benefiting from the advanced knowledge in question.
We called the rule in question the advanced knowledge rule. Note that knowing in advance one of the possible halves of the information about the solution one will produce and measure in the future and benefiting from this knowledge to produce the solution with fewer computation steps is a causal loop. We can therefore say that, according to this rule, the quantum speedup is explained by a mutually exclusive or of causal loops, one for each possible way of taking half of the information about the solution.
By the way, here and in the following by quantum speedup we will always mean its certain form, the fact that the quantum algorithm needs fewer computation steps than classically possible. Of course, this is not the case of most quantum algorithms, which need fewer computation steps than their best classical counterpart, but whether the latter is the best classically possible is not known.
We should also note that the teleological character of quantum algorithms—their being evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again)—is implicit in the rule in question. Also the notion of finalism acquires a precise definition. The teleological evolution would not be attracted by some scientifically undefinable goal but, very simply, by its own end product (in quantum algorithms, the solution of the problem).
Until now, this advanced knowledge rule was, so to speak, an empirical rule shown to be satisfied by all major quantum algorithms. In the present work, we should have demonstrated it, incidentally, in a surprisingly simple way. Of course, we would have consequently demonstrated the teleological character of quantum algorithms.
As a matter of fact, we have provided two redundant demonstrations of this rule, what should not be despised given its destabilizing character. Let us outline them.
One is an application of classical logic to what the quantum algorithms does: providing a quantum speedup with respect to what is classically possible.
By the way, we will have to use a quantum description of the quantum algorithms that is not their customary quantum description. The latter is limited to the unitary process of solving the problem, where the quantum speedup resides, and the final measurement of the solution. It is therefore incomplete; the canonically complete description of a quantum process consists of initial measurement, unitary evolution, and final measurement. We complete the customary quantum description by adding to the description of the process of solving the problem that of the upstream process of setting the problem. The result is: an initial measurement of the problem setting in a quantum superposition (or, indifferently, mixture) of all the possible problem-settings selects one of them at random; the unitary computation and measurement of the corresponding solution follow. Note that this quantum description only involves the measurements of two commuting observables (the problem-setting and the solution of the problem); therefore, it should be subject to classical logic. As we shall see in the next section, the classical logic description of the quantum speedup is indeed the advanced knowledge rule.
Thus, in particular, this classical logic description of the quantum algorithm would tell more than its present quantum description; we should conclude that the latter is still incomplete. Indeed, it is simply completed by time-symmetrizing it. We shall see that the time-symmetrization of the quantum description yields the quantum superposition of the causal loops involved in the advanced knowledge rule. In other words, the time-symmetrization of the quantum description of the quantum algorithm generates the physical correspondent of the advanced knowledge rule.
The second demonstration of the advanced knowledge rule simply comes from the fact that this time-symmetrization of the quantum description of the quantum algorithm turns out to be mandatory for another reason independent from the one seen so far. As we shall see, it is to satisfy an atemporal symmetry implicit in the quantum description. Of course, this time-symmetrization yields the physical correspondent of the advanced knowledge rule—it is a demonstration of the rule.
Of course, either demonstration of the advanced knowledge rule is also a demonstration of the teleological character of quantum algorithms, their being evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again).
A further step, important for what concerns the possible teleological character of the evolutions of the living, is that all the above generalizes to quantum correlation. For simplicity, we limit ourselves to the case that, in the quantum algorithm, the solution is an invertible function of the problem-setting and is determined with certainty. We can then say that the quantum algorithm builds up one-to-one correlation between two initially unrelated observables (the problem-setting and the solution of the problem) without ever changing the eigenvalue of the observable selected by the initial measurement (the problem-setting). We shall show that any building up of quantum correlation of this kind (not only that in quantum algorithms) can occur with quantum speedup and therefore be teleological in character. Note that the “can be” means that there is certainly room for achieving a quantum speedup, but, of course, whether it is achieved, in full or in part or not at all, depends on the character of the unitary evolution that builds up the quantum correlation. By the way, in quantum search algorithms, achieving the quantum speedup is not difficult: [7] shows that even a random quantum search for the solution yields a just suboptimal quantum speedup.
With this, we can go to the evolutions of the living for which the teleological notion was originally conceived.
First, let us note that the dismissal of the teleological notion was not without (scientific) regret; indeed, it meant doing without a possibly enlightening relation between the notion of evolution of life and that of the functional structure of the living. We cite two authoritative objections to this dismissal. One is a famous sentence by the evolutionary biologist Haldane [8]:
Teleology is like mistress to a biologist, he cannot live without her, but is unwilling to be seen with her in public.
The other is the position of the natural philosopher Nagel [9], who provides a series of very practical arguments for which the mechanistic, neo-Darwinian conception of nature could not account for the fundamental features of the living, and this in favor of a holistic and teleological view. In our possibly biased opinion, this would be a case in which common sense prevails over a de facto no-retrocausality precept that, according to the present work, would be a prejudice.
Let us now go to what the teleological character of quantum algorithms can say to the evolutions of the living.
Naturally, the first observation must be on the enormous difference between quantum algorithms and the evolutions of the living. The former have maximum durations of a few dozen microseconds and a number of quantum bits limited to the order of one thousand; beyond that, in today’s implementations, there is quantum decoherence. The latter have a duration of billions of years and can have the size of the Earth. However, interestingly, the very context we are dealing with suggests a way out. The difference is enormous in the usual classical way of viewing the macroscopic reality, but disappears under the quantum cosmological assumption that the universe evolves in a unitary quantum way [10, 11]. Furthermore, surprisingly, under the quantum cosmological assumption and for the Fine-tuned Universe version of the Anthropic Principle [11], the evolution of the universe toward the “value of life” (in particular“existence of sentient life” and “absence of life”) acquires the form of a quantum algorithm and, as such, it can occur with quantum speedup and thus be teleological in character.
Let us describe how things should go.
The Fine-tuned universe version of the Anthropic Principle is about the very relevant fact that the least change of the value that the fundamental physical constants have in the universe we are in would have developed a relatively trivial universe incapable of producing life.
Under the quantum cosmological assumption, this version of the Principle can be formalized as follows. We should introduce two commuting observables of the same size: the value of the fundamental constants and the “value of life”; the latter should assume in particular the two values: “existence of sentient life” and “absence of life”.
Now, always under the quantum cosmological assumption, we should consider a quantum superposition of all the possible universes, all with the same fundamental physical laws (thus, the same unitary transformation from the beginning of time until today) but for the value of the fundamental physical constants, and each with its own value of the latter. Initially, the value of the fundamental constants and that of life are two independent observables: indeed, no matter the value of the fundamental constants, the value of life is “absence of life”. The subsequent unitary evolution of this quantum superposition, from the beginning of time until today, builds up a one-to-one correlation between the value of the fundamental constants and the value of life. Today, in the universe we are in, the value of the fundamental constants is that which produces sentient life; in the other universes, where this value is different, there is absence of life (theory says that, the absence of life in the other universes today is already definitive).
For what we have said about the generalization of the quantum speedup from quantum algorithms to the building up of quantum correlation, the evolution of the universe toward the value of life can occur with quantum speedup and, therefore, be teleological in character. The “can occur” is in the sense that there is room for its occurring with quantum speedup.
Of course, in the universe we are in, the possible quantum speedup and the consequent teleological character is for the evolution of the universe toward life. We will see that this possibility would be inherited by the evolution of life on Earth, in principle, the trace over all the rest of the evolution of the universe toward life. Then, for consistency, it would be inherited by the evolutions of species.
Summing up, while the teleological character of quantum algorithms, as we shall see, should be certain, the fact that—under the quantum cosmological assumption—it would extend to the evolutions of the living should be considered as plausible as the quantum cosmological assumption. However, note that all this is quite different from the present widespread belief that the notion of teleological evolution of the living is an absurdity.
Let us eventually note that, so far, holism and finalism have been the enemy of the reductionist and mechanistic vision of science. In the present theory of the quantum computational speedup, they would acquire a physical basis. Of course, this would not touch on the immense practical value of the mechanistic and reductionist view. That the description of the evolution of any system is the trace over all the rest of the quantum description of the evolution of the universe obviously holds only in principle and does not give what the classical description gives. However, this might pinpoint a direction of the research. At the theoretical level, it could be interesting to study how this “tracing over all the rest” conciliates a holistic and teleological view in the longer term with a mechanistic and reductionist view in the shorter term. This might even be a further perspective under which to explore the unification between classical and quantum mechanics.
2 Time-symmetric explanation of the quantum speedup
We provide a description of quantum algorithms and their quantum speedups open to the coexistence of causality and retrocausality. Since this is a way of thinking quantum computation radically different from the usual one, we provide it first at a conceptual level. All formalizations are postponed in the Appendix. The section is divided into subsections corresponding to the various arguments.
Let us first acknowledge the fact that that of the quantum computational speedup has been one of the fundamental discoveries of the past century. Together with that of quantum communication, it has created an entirely new branch of physics. This said, for the sake of clarity, we will be completely frank in criticizing what we think are the shortcomings of the current way of viewing quantum computation.
2.1 On the explanation of the quantum computational speedup
We repeat here for convenience the advanced knowledge rule.
In quantum algorithms, it is as if the problem-solver always knew in advance, before beginning her problem-solving action, any of the possible halves of the information that specifies the solution of the problem she will produce and measure in the future and could use this knowledge to produce the solution with fewer computation steps. The number of computation steps needed in the optimal quantum case is that needed in the optimal logical case benefiting from the advanced knowledge in question.
Of course, this rule would offer a unified quantitative explanation of the quantum speedups that is still lacking in mainstream quantum information. Let us discuss this point. To the best of our knowledge, the absence of a unified quantitative explanation has never been acknowledged by the scientific community.
People currently say that there is an explanation of the quantum speedup; it would be that of quantum parallel computation, the fact that many computations are performed at the same time in quantum superposition. Indeed, this captures the essential character of quantum computation, but, being a qualitative explanation of the quantum speedup, which is a quantitative feature, cannot be considered its explanation. By the way, what distinguishes science from opinion is, in particular, its quantitative character. The fact that the most popular explanation of the quantum speedup, a quantitative feature, is qualitative in character, would seem to be rather unscientific. The only justification we can see is that this is a way of removing the fact that, forty years after its discovery—an explanation of the quantum speedup is still missing. This is naturally our opinion. However, as a matter of fact, outside our (unconventional) retrocasusal approach to the quantum speedup, in the literature, there is no trace whatsoever of any search for a unifying-quantitative explanation of it.
Of course, as already noted, if the only way of providing the unified quantitative explanation in question is by resorting to retrocausality and finalism, the de facto rejection of these notions by the scientific community might have played a role.
Neither can be considered an explanation the fact that each quantum algorithm naturally explains its quantum speedup. In fact, it would be an explanation completely different for different quantum algorithms. It would be like having different explanations for different throws of a stone without knowing that they all describe a parabola. Obviously, unification is a fundamental objective of any description of physical phenomena.
For what concerns the current mainstream explanation of the quantum speedup, let us say, it is the case of crying out that the emperor is naked. This emphasis is to underline the absence of any recognition of the fact that the current view of quantum computation has a shortcoming.
Of course, the advanced knowledge rule would exactly be that unifying-quantitative explanation. However, to date, it has had a twofold problem: first, although shown to apply to all the major quantum algorithms, it was not demonstrated; second, it goes against the deeply rooted de facto ban on retrocausality. In this work, we should have demonstrated it, moreover, in a surprisingly simple way. As a matter of fact, we shall be dealing with two independent and redundant demonstrations, which is not to be despised because of the unconventional character of the rule.
2.2 First demonstration of the advanced knoledge rule
We have already given in the Inroduction the quantum description of quantum algorithms extended to the process of solving the problem. An initial measurement in a quantum superposition (or, indifferently, mixture) of all the possible problem-settings selects one of them at random [^1] The problem-setter could unitarily change it at will; this, here, would change nothing.. The unitary computation and the final measurement of the corresponding solution follow. For the sake of simplicity, we consider the case that the solution is an invertible function of the problem-setting and is determined with certainty.
Now, it should also be noted that this quantum description of the quantum algorithm is with respect to the customary observer external to the quantum process. To present ends, we should also relativize it with respect to the problem-solver, who computes the solution of the problem. Of course, the advanced knowledge of one of the possible halves of the information about the solution we are dealing with will have to appear in the quantum description of the quantum algorithm with respect to her. Now, from her, the problem-setting selected by the initial measurement should be concealed: as we shall see in the following, it would also tell her the solution of the problem before she begins her problem-solving action. Since the unitary computation of the solution never changes the problem-setting selected by the initial measurement, this concealment is easily obtained. It suffices to postpone the projection of the quantum state associated with the initial measurement to the end of the unitary evolution.
This concealment is obtained by postponing to the end of her unitary computation of the solution the projection of the quantum state associated with the initial measurement; see Section (2) of the Appendix.
Let us call this extended quantum description with respect to the problem-solver “the ordinary quantum description of the quantum algorithm”. It is ordinary because it is still the customary “causal” quantum description of mainstream quantum mechanics—where causality always goes in the same time-direction.
We can go now to the first demonstration of the advanced knowledge rule. It simply consists in describing in a classical logic way what the quantum algorithm does—computing the solution of the problem with a number of computation steps that is less than the minimum possible in the classical case.
By the way, involving only the measurements of commuting observables (the problem-setting and the solution of the problem) quantum algorithms, and what they do, should be subject to classical logic.
To present ends, we need a quantum algorithm that certainly gives a quantum speedup with respect to what is classically possible. Now, most quantum algorithms yield a quantum speedup with respect to their best classical counterpart, but whether the latter is the best possible is not known. Excluding elementary one-computation-step quantum algorithms whose dimensions do not scale up, we find our choice limited to Grover’s quantum search algorithm [12]. Besides yielding a well determined quantum speedup with respect to what is classically possible, it has the following convenient features: (i) It is a most universal quantum algorithm that can solve any NP problem with a quadratic quantum speedup; more than a single quantum algorithm, it is an entire test bed on which to test the present theory of the quantum speedup. (ii) In Long’s [13] version, it always provides the solution with certainty; this simplifies things. (iii) In it, the solution is an invertible function of the problem-setting; this simplifies its time-symmetrization. (iv) It is optimal in character; this, of course, constitutes an important reference. By the way, note that the notion of searching for something and finding it is most universal and this universality is completely represented by Grover algorithm.
We show that the classical logic description of the quantum speedup delivered by Grover algorithm implies the advanced knowledge rule. We begin with the simplest instance of Grover algorithm. As we shall soon see, everything immediately generalizes first to the full fledged Grover algorithm and then to most general forms of quantum correlation (in quantum algorithms, that between the problem-setting and the solution of the problem).
The problem to be solved is as follows. Bob, the problem-setter, hides a ball in a chest of four drawers. Alice, the problem-solver, is to locate the ball by opening drawers. In the worst classical case, she has to open three drawers. If the ball is only in the third drawer opened, she has solved the problem. If not, it must be in the only drawer not yet opened, and she has solved the problem of locating the ball as well.
Instead, Grover algorithm (its simplest instance) always solves the problem by opening just one drawer, in a quantum superposition of the four possible drawer numbers. Obviously, always solving the problem by opening just one drawer would be impossible in the classical case: indeed, we are dealing with a quantum computational speedup.
There is a classical logic equivalent to the fact that the quantum algorithm always solves the problem by opening a single drawer. By the way, since the fact that a single drawer is opened is logically true in itself, we can forget the specification that this is done for a quantum superposition of the four drawer numbers.
The logical equivalence is as follows. Logically, it is as if the problem-solver always knew in advance, before beginning her problem-solving action, that the ball is in a given pair of drawers (one of which, of course, the one with the ball) and took advantage of this knowledge to solve the problem by opening just one of them. In fact, if the ball is in the opened drawer, she has solved the problem; if not, it must be in the other drawer and she has solved the problem of locating the ball as well, in either case having opened just one drawer. This also means knowing in advance half of the information that specifies the solution of the problem that will be produced and measured in the future.
Let us open a parentheses to observe that we are not talking about the quantum algorithm as it were a classical algorithm. In other words, here, classical logic should not be confused with classical physics, which might be a temptation in the present case.
Always knowing in advance that the ball is in a given pair of drawers, in a mutually exclusive way, one of the possible pairs that host the drawer with the ball, is logically necessary to always solve the problem by opening just one drawer. Then the question becomes: from where does Alice, before beginning her problem-solving action, always get that information about the solution? Since, for her, the only possible source of it is the solution she will produce and measure in the future, the only possible answer is that the information in question retrocausally comes to her from her future measurement of the solution. This is formalized in Section (3) of the Appendix. However, before looking at it, one should read the description of the time-symmetrization of the quantum algorithm that we shall provide in a moment.
Note that we are dealing with a causal loop. Alice knows in advance one of the possible halves of the information about the solution she will produce and measure in the future and can use this knowledge to produce the solution with fewer computation steps (drawer openings); the fewest is the number needed in the optimal logical case benefiting of half of the information about the solution. There is a causal loop for each way of taking half of the information about the solution. Therefore, we can say that this quantum speedup is logically explained by the mutually exclusive or of the causal loops we are dealing with. Of course, this is the advanced knowledge rule for the present instance of Grover algorithm, logically implied by the sheer existence of the quantum speedup. In summary, the classical logic description of the quantum speedup delivered by (the simplest instance of) Grover algorithm implies the advanced knowledge rule.
This explanation of the quantum speedup immediately generalizes to any number $N$ of drawers. As is well known, the number of drawer openings needed by the (optimal) Grover algorithm is on the order of $\sqrt{N}$ against the order of $N$ of the best classical case, hence its quadratic quantum speedup. This is logically equivalent to knowing in advance, in a mutually exclusive way, that the ball is in one of the possible tuples of $\sqrt{N}$ drawers and benefiting from this knowledge to locate the ball in the best logical way by opening the order of $\sqrt{N}$ drawers. Of course, this is what the optimal Grover algorithm does. Also in this case, using the advanced knowledge of half of the information about the solution in an optimal logical way yields the optimal quantum speedup of Grover algorithm.
Note that, in equivalent terms, the computational complexity of the problem to be solved reduces, from searching the ball in $N$ drawers to searching it in $\sqrt{N}$ drawers. In still equivalent terms, the dimension of the Hilbert space in which the search for the solution takes place reduces from $N$ to $\sqrt{N}$.
Holding for Grover algorithm, the above retrocausal explanation of the quantum speedup must hold for the process of solving with a quadratic quantum speedup any NP problem. Here we do not care about the fact that most NP problems are classically solved with a number of computation steps exponential in their size, so that a quadratic quantum speedup can be seen as a little advantage. Of course, for the present ends, we are just interested in having a quantum speedup with respect to what is classically possible.
We shall see that there are further generalizations in the following of this section. In view of them, note that, in order to derive the present explanation of the quantum speedup, we never needed to know how Grover algorithm achieves the latter, we only needed to know that it achieves it.
2.3 Completing the ordinary quantum description of the quantum algorithms by time-symmetrizing it
The above subsection also says that the ordinary quantum description of quantum algorithms, which obviously does not describe the advanced knowledge rule, is incomplete. Of course, to bring it in line with the classical logic description, it should be completed. It turns out that this is simply done by time-symmetrizing it.
We describe this operation. For the sake of simplicity, we consider the case that the solution is determined with certainty and is an invertible function of the problem-setting—exactly as in Grover algorithm. We should evenly share between the initial measurement of the problem-setting and the final measurement of the solution the selection of the information that specifies both; of course, in the ordinary quantum description of the quantum algorithm, all the information is selected by the initial measurement. Naturally, there is a plurality of possible ways of evenly sharing: we should take their quantum superposition.
In each time-symmetrization instance, the selection performed by the initial measurement propagates forward in time by the unitary computation of the solution until it becomes the state immediately before the final measurement. Time-symmetrically, that performed by the final measurement propagates backward in time by the inverse of the unitary computation of the solution until replacing the previous outcome of the initial measurement. This backward in time propagation, which inherits both selections, is an instance of the time-symmetrized quantum algorithm. As we can see in Section (3) of the Appendix, in the initial measurement outcome of the time-symmetrized instance of the quantum algorithm it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future; she can then use this knowledge to compute the solution with fewer computation steps. This is the physical correspondent of one of the logical causal loops whose mutually exclusive or is logically implied by the advanced knowledge rule.
Eventually, we should take the quantum superposition of all the time-symmetrization instances. The time-symmetrized quantum description of the quantum algorithm (i.e. its completed quantum description) becomes the quantum superposition of the causal loops whose mutually exclusive or is logically implied by the quantum speedup: So completed, also the quantum description of the quantum algorithm describes the advanced knowledge rule described by its classical logic description—it describes its physical counterpart.
2.4 Second demonstration of the advanced knowledge rule
The second demonstration of the advanced knowledge rule simply comes from the observation that the time-symmetrization of the ordinary quantum description of the quantum algorithm, which provides the (physical counterpart of the) advanced knowledge rule, is mandatory also for another reason, as follows.
In the ordinary quantum description of the quantum algorithm, of course, the unitary computation of the solution never changes the corresponding problem-setting. Therefore, one should be free to postpone the initial measurement of the problem-setting along the subsequent unitary computation of the solution until it is simultaneous with the final measurement of the solution. In this case, we find ourselves with the simultaneous measurements of two maximally entangled observables (the problem-setting and the solution, one an invertible function of the other), and the time-symmetrization we are dealing with must already be in place to represent the perfect (atemporal) symmetry between these two measurements. Of course, there would be no reason why a measurement selected more information than the other. As we have just seen, this time symmetrization generates the quantum superposition of the causal loops whose mutually exclusive or is logically implied by the sheer existence of the quantum speedup—the physical counterpart of the advanced knowledge rule. This is indeed the second demonstration of the advanced knowledge rule.
By the way, this result is reminiscent of Wheeler and Feynman’s [14] time-symmetric absorber theory; we would say: with the advantage that the time-symmetrization we are presently dealing with relies on a reversibility of the quantum process that, unlike the classical one, is such in both an ideal and objective way.
2.5 Origin of the retrocausality
A quantum algorithm is a relatively complex process. A natural question is whether the retrocausality implicit in the classical logic description of its overall result, i.e. the quantum speedup, derives from some more elementary part of it. Of course, in its ordinary quantum description, the unitary computation of the solution is an alternation of generations of quantum state superpositions and quantum interferences. Since the latter have a classical analog, the essentially quantum generation of a quantum state superposition is the suspect.
Let us first provide a very intuitive exposition of the fact that the classical logic description of a state superposition necessarily involves the coexistence of causality and retrocausality. We resort to a prescientific notion of state superposition. It is the legend of the monk seen by his disciples on the two banks of a river at the same time: we should provide a classical logic description of this. Being on one and the other bank is possible in two different times but, at the same time, from the ordinary classical logic standpoint, they should be two mutually exclusive possibilities. Until now, being on the two banks at the same time would seem to be a logical impossibility. However, let us note that it would be impossible in a classical logic description closed to retrocausality. Opening classical logic to the coexistence of causality and retrocausality completely changes the situation [^2] By the way, classical logic is neutral for what concerns the notion of causality. Customarily associating it with a causality that always goes in the same time-direction must be consequent to the de facto no-retrocausality precept. In principle, classical logic should also go along with the coexistence of causality and retrocausality.. Indeed, the monk could have gone to one of the two banks along a certain path, then (then in causal order) he could have gone backward in time undoing the path just followed (like in a reversed film); then, going forward in time again, he could have reached the other bank along a different path. This causality zigzag, of course, makes its being on the two banks at the same time logically possible.
Let us now bring the above classical logic description of a state superposition to a quantum state superposition. Think of a photon going in quantum superposition through the two branches of a Mach-Zehnder interferometer. The two branches should be tuned in such a way that the photon always ends up in the same detector, so that we have a (of course, time-reversible) unitary evolution that goes from before the photon enters the interferometer to immediately before its detection.
The classical logic description of this can be as follows. In entering the interferometer, the photon takes, say, the upper branch until it reaches the state immediately before detection (by the detector in question). Of course, this also provides the phase of the photon along the path followed. Then (in causal order), the photon goes backward in time undoing the path just followed until before entering the interferometer. Then, it goes forward in time again taking the lower branch until reaching the detector in question and, this time, being detected. The zigzag path followed in this way by the photon through the interferometer, with the associated photon phases, of course, describes in a classical logic way the fact that the photon goes through the interferometer in a quantum superposition of the upper and lower branches, eventually ending up in that detector. It should be clear that, also in this case, the involvement of retrocausality in the classical logic description is necessary.
We should think that the classical logic description of the sequence of generations of quantum state superpositions and interferences that constitutes the unitary computation of the solution, in the end, simplifies into a description that implies the advanced knowledge rule.
Summarizing, at the root of the logical form of retrocausality we are dealing with there should be the fact that the classical logic description of the generation of a quantum state superposition (one of the pillars of quantum mechanics) necessarily involves it. By the way, note that this is not an idle speculation, as might even seem. Indeed, we are dealing with a logical form of retrocausality with huge consequences, such as providing a physical basis to the notion of teleological evolution.
2.6 Generalization
In the second demonstration of the advanced knowledge rule, we have shown that a unitary evolution between two one-to-one correlated measurement outcomes that never changes the eigenvalue selected by the initial measurement is incomplete and must be completed by time-symmetrizing it. Note that we never needed to look inside that unitary evolution $\hat{\mathbf{U}}_{1,2}$: see Tables I through IV of the Appendix. Therefore, this time-symmetrization does not only hold for quantum algorithms, it lends itself to an important generalization:
The ordinary quantum description of any unitary evolution between two one-to-one correlated measurement outcomes that does not change the eigenvalue selected by the initial measurement is incomplete and is completed by time-symmetrizing it.
Let us call it the quantum correlation time-symmetrization rule.
The case where the one-to-one correlation is built up by the unitary evolution characterizes the quantum search algorithm, where this building up occurs with quadratic quantum speedup. However, the possibility of occurring with quadratic quantum speedup holds for any unitary evolution that responds to the rule in question and builds up one-to-one quantum correlation between two initially unrelated observables. Indeed, that this process can occur with quadratic quantum speedup is what the time-symmetrized quantum description of the quantum process says—see the Appendix. Of course, in the case where we are not dealing with a quantum algorithm, the notion of advanced knowledge of one of the possible halves of the information about the solution must be replaced by the equivalent notion of quadratic reduction of the dimension of the Hilbert space in which the search for that one-to-one correlation takes place, from $N$ to $\sqrt{N}$.
The case in which the one-to-one quantum correlation is not built up by the unitary evolution because it preexists it (more precisely, what would preexist is the maximum entanglement between the observables of the two measurements) is, for example, that of quantum nonlocality. Here, we have a unitary evolution that spatially separates the subsystems of two already maximally entangled observables. With the two measurements performed in the same basis and the spatial separation process preceding the first measurement, this responds to the quantum correlation time-symmetrization rule. The time-symmetrization of the ordinary quantum description of this form of quantum nonlocality yields a quantum superposition of causal loops that locally explain it and play the role of the hidden variables envisaged by Einstein et al [15]. Indeed, these causal loops locally go from one to the other measurement outcome via the time the two subsystems were not yet space separate, where they locally tell the subsystem of the not yet measured observable the state to assume in view of the future measurement—see [5]. Note that this is what Einstein et al required from their hidden variables. That was, of course, without resorting to retrocausality, but under a deterministic view of the measurement process.
In summary, this application of the quantum correlation time-symmetrization rule both demonstrates the incompleteness of the ordinary quantum description of quantum nonlocality and provides the hidden variables that complete it. Of course, this also means that the incompleteness we are presently ascribing to the ordinary quantum description of quantum algorithms is the same that Einstein and others ascribed to the quantum description of quantum nonlocality.
By the way, the fact that the present explanation of the quantum speedup also demonstrates the incompleteness of the quantum description highlighted by Einstein et al in their EPR paper [15] might be destabilizing for the following reason: After highlighting for the first time quantum nonlocality and inspiring fundamental works like Bell’s theorem [16] and Ekert’s quantum cryptography based on it [17], their argument had been dismissed by the physicists’ community because of the alleged absence of a satisfactory way of completing the quantum description by means of local hidden variables.
2.7 About the teleological character of the evolutions of the living
We have already seen in the Introduction that, under the quantum cosmological assumption and for the Fine-tuned Universe version of the Anthropic Principle, the physical basis of the teleological character of quantum algorithms also applies to the evolution of the universe toward the value of life. We repeat here for convenience the reason for its possible quantum speedup and consequent teleological character.
Under the quantum cosmological assumption, we should consider a quantum superposition of all the possible universes, all with the same fundamental laws (thus, the same unitary transformation from the beginning of time until today) but for the value of the fundamental constants, and each with its own value of the latter. Initially, the value of the fundamental constants and that of life are two independent observables: indeed, no matter the value of the fundamental constants, the value of life is “absence of life”. The subsequent unitary evolution of this quantum superposition, from the beginning of time until today, builds up a one-to-one correlation between the value of the fundamental constants and the value of life. Today, in the universe we are in, the value of the fundamental constants is that which produces sentient life, in the other universes where this value is different there is absence of life.
We have seen in the Generalization subsection that any such building up of quantum correlation can always occur with quantum speedup in the sense that there is certainly the room for it. It can consequently be teleological in character.
Of course, in the universe we are in, the possible teleological character of the evolution of the universe toward the value of life holds for the evolution of the universe toward life. Then, it is easily seen that this possible teleological character must be inherited by the evolution of life on Earth. In principle, the latter evolution can be derived from the former evolution by tracing all the rest. Therefore, we can speak of an image of the evolution of life on Earth in the evolution of the universe toward life. Then, first, evolution times must be the same for the evolution of life on Earth and its image in the evolution of the universe. Second, the amount of physical resources used by the evolution of life on Earth cannot exceed the amount of physical resources used by its image in the evolution of the universe. In fact, the latter resources comprise the former.
In summary, under the quantum cosmological assumption, the evolution of life on Earth can occur with quantum speedup and consequently be teleological in character; it might actually occur with quantum speedup if the search for the value of life by the evolution of the universe toward life were random. Then, just for consistency, the evolutions of species would also be the same.
Not to hide behind a finger, let us exemplify what a teleological character of the evolution of life would mean for the notion of the evolutions of species. It is known that, for many millions of years, the ancestors of the birds were feathered dinosaurs that did not fly. Very light, strong but flexible, cave and air-proof, even at that time feathers were ideal for flight, very close to those of the future birds. So why anticipate something designed for flight many millions of years in advance?
The usual explanation, of course, is causal. First, an ancestor of feathers appeared by random DNA mutation, then evolved as a good thermal insulator and perhaps to speed escape from predators; only eventually, feathers turned out to be what was needed to conquer the air.
All this could remain true in the teleological view with just one, but very significant, addition: Those random mutations that would have led in the future to the conquest of the air would have more likely occurred. See our preprint [6] for a more comprehensive discussion of the teleological notion.
All this is obviously destabilizing; it is a way of viewing the macroscopic reality that goes against the current (scientific) way of viewing it, and revamps the natural philosophy views that were ascribed by science to the unscientific beliefs. Of course, the fact that what occurs to us today is affected by the drying up of an area of Africa three millions of years ago is obvious, but that it is also affected by what we will be in three million years’ time, to the current way of thinking is more difficult to digest.
This also raises the question of how to describe a macroscopic evolution endowed with quantum speedup. Of course, classical physics would turn out to be unable to describe not only the processes of the microscopic world but also the teleological evolutions of the macroscopic world.
We quote a precedent in the quantum description of macroscopic natural evolutions. It is about the asymptotic behavior of predator species competing for the same prey [18]. It is explained by a quantum model, whereas it would be difficult to explain it in a classical way. In our case, the canonical way of describing the evolution of life on Earth would be to obtain it from the description of the evolution of the universe toward life by tracing-out all the rest. Since this obviously holds only in principle, we should also model the macroscopic teleological evolutions of species endowed with quantum speedup in a quantum way. In other words, we should accept quantum superposition, entanglement, and correlation between the observables of macroscopic objects.
This might even be a good approximation, since, in the universe we are in, the evolution of life on Earth in its time lapse should have been almost autonomous from the rest of the unitary quantum evolution of the universe toward life. Indeed, if it were completely autonomous, of course, it would be a unitary quantum evolution itself.
Incidentally, it might also be interesting to derive the description of the evolution of life on Earth from that of the universe toward life by tracing out all the rest at the level of toy models of the two evolutions. Within the present theory, this might highlight significant characters of a possible teleological evolution of life on Earth.
By the way, of course, the quantum cosmological view of the evolution of the universe would “unify” quantum and classical physics simply by canceling the latter. However, naturally, classical physics turns out to work extremely well in describing macroscopic reality. Presently, it would have the shortcoming of being unable to describe macroscopic teleological evolutions. This situation might point to a particular object of research. It could be interesting to study how the operation of “tracing-out all the rest” (at the toy model level) reconciles the teleological character of an evolution in the longer term with its classical character in the shorter term. This could also be a further perspective under which to address the problem of unifying classical and quantum theory.
Let us end the present section by recalling what, to our knowledge, are the key steps that led to the discovery of the quantum computational speedup. According to the present theory, this discovery would be even more fundamental than currently thought. The first step was showing that computation is possible in the quantum world and introducing the fundamental notion of quantum bit, by Finkelstein [19] in 1969; that was also the first time that quantum physics moved to the world of the complex, indeed to a cellular automaton model of the universe—this was in the context of an attempt to unify relativity and quantum theory. Then, there was the introduction of the notion of logically reversible computation by Bennett [20] in 1973, of course, a prerequisite for the building up of computations in unitary quantum evolutions. The discovery of the superior computational efficiency of the quantum world with respect to that of the classical world in 1982 is due to Feynman [21]. Still in 1982, there has been the introduction of the notion of conservative logic by Fredkin and Toffoli [22]: with it, we have the logically reversible gates needed to perform quantum computations. The seminal quantum algorithm that delivers a quantum speedup was conceived by Deutsch [23] in 1985.
Note how the notion of the logical reversibility of computation played an essential role in the development of the notion of quantum computation. The present time-symmetric quantum description of quantum algorithms could be a further step in this direction. Involving a time-symmetric coexistence of the two possible time-directions of causality, it could be said that it exploits the objectively ideal time-reversibility of unitary quantum evolutions in the possibly deepest way.
3 Precedents
We have seen that applying the quantum correlation time-symmetrization rule to quantum nonlocality provides a local explanation of it. Now, we recall a historical precedent of this.
It is a work by Costa de Beauregard [24] that also provides a retrocausal explanation of quantum nonlocality. In it, the outcome of the first performed measurement—of the observable of one of the two space separate subsystems—locally and unitarily goes backward in time until the time the two subsystems were not yet space-separate, where it locally tells the other subsystem the state to assume in view of the future measurement. Naturally, we are not dealing with the present time-symmetric form of retrocausality, but the concept is similar. Interestingly, at the time, Costa de Beauregard was working for de Broglie, who forbid him to publish that work involving retrocausality for years. Ultimately, that ban was withdrawn only after the publication of Wheeler and Feynman’s work on the time-symmetric emitter-absorber theory, which also implied a form of retrocausality. See [25] for an account of this episode. It is rather significant of the attitude, of even top physicists, toward the retrocausality notion.
Speaking of precedents, of course, the present work is inspired by time-symmetric quantum mechanics [26, 27]. Furthermore, there have been and there are studies on the possible violations of the causal principle by quantum mechanics. However, first, none of them concerns the quantum computational speedup and, second, they all apply to very specific physical situations to which they might be confined. On the contrary, the present time-symmetric form of retrocausality is sort of universal for quantum correlations. The possible relation of those works to the present work, as things are now, would not be evident. Of course, it could be the object of study.
4 Conclusions
The present work goes against a long-standing de facto no-retrocausality tradition of physics. More specifically, since causality could well go backward in time along a unitary quantum evolution that is both objectively and ideally time-reversible, we should say that, in the quantum world, the scientific tradition amounts to asserting the mutual exclusivity between the two possible time directions of causality.
Let us first address the reliability of the work, a particular issue because of its unconventionality.
That the classical logic description of the quantum speedup must resort to retrocausality should be obvious. Indeed, it should be obvious the fact that always locating a ball hidden in one of four drawers by opening a single drawer is logically equivalent to always knowing in advance that the ball is in a pair of drawers. To the problem-solver, the only possible source of this information is the solution of the problem she will produce and measure in the future: hence retrocausality. We have seen how this immediately generalizes to the full fledged Grover algorithm and a most general form of quantum correlation.
At this point, we should address the legitimacy of applying classical logic to what the quantum algorithm does. Let us recall our two arguments in favor of this choice.
The first is that, only involving the measurements of commuting observables (the problem-setting and the solution of the problem), quantum algorithms (and what they do) should be subject to classical logic.
The other is that all the results obtained under the assumption that quantum algorithms can be described by classical logic are also obtained by time-symmetrizing their ordinary quantum description; this, in turn, is mandatory to satisfy the atemporal symmetry between the simultaneous measurements of two maximally entangled observables implicit in the ordinary quantum description of quantum algorithms.
We would think that the present retrocausal explanation of the quantum speedup is reliable enough to deserve a discussion despite its destabilizing consequences (or, according to tastes, even right for them).
The other side of the coin of the destabilizing character of the work is, of course, its revolutionary character, always due to the retrocausality implicit in it. We should distinguish between what should be certain and what could be just plausible. Beginning with what should be the certain: (i) The work provides a unified quantitative explanation of the quantum speedup, which, forty years after its discovery, was still missing from mainstream quantum mechanics. (ii) It theoretically confirms the incompleteness of the quantum description of quantum nonlocality sustained by Einstein and others, also providing the local hidden variables envisaged by those authors to complete it. (iii) It highlights the teleological character of quantum algorithms; by the way, given the negative view that the teleological notion has today in science and the scope of this notion, this should be the most revolutionary result of all. What is just plausible: It shows that, under a quantum cosmological interpretation of the Fine-tuned Universe version of the Anthropic Principle, the physical basis of the teleological character of quantum algorithms applies as well to the evolutions of the living.
Let us eventually note that all this would change our way of viewing physical reality. Curiously, it would be a case where the progress of science, the advent of quantum mechanics, brought us back to a way of viewing physical reality that preceded science.
Appendix
We have seen that, according to the present theory, the usual quantum description of quantum algorithms is incomplete. How to complete it has already been discussed at a conceptual level in the previous part of the work. We now provide the formalization. We do it in the simplest instance of Grover algorithm; we shall see that everything immediately generalizes.
We should perform the following steps:
(1) Extend the usual quantum description of the quantum algorithm, limited to the process of solving the problem, to the upstream process of setting the problem: One should measure the problem-setting in a quantum superposition of all the possible problem-settings. The unitary computation and measurement of the corresponding solution follow.
(2) Relativize the extended quantum description of the quantum algorithm with respect to the problem-solver, from whom the problem-setting selected by the initial measurement should be concealed (it would tell her also the solution of the problem, before she begins her problem-solving action). Note that the advanced knowledge of one of the possible halves of the information about the solution of the problem must be in the quantum description of the quantum algorithm with respect to the problem-solver.
(3) Complete the quantum description of the quantum algorithm obtained so far by time-symmetrizing it.
Steps (1) through (3) are developed in the following correspondingly numbered sections.
(1) Extending the usual quantum description of quantum algorithms to the process of setting the problem
The usual quantum description of quantum algorithms is limited to the unitary process of computing the solution of the problem and the final measurement of the solution. We extend it to the upstream process of setting the problem by measuring the problem-setting in a quantum superposition of all the possible problem settings.
We introduce the notation; we make reference to the simplest, four drawer, instance of Grover algorithm:
We number the four drawers in binary notation: $00,01,10,11$.
We will need two quantum registers of two quantum bits each:
A register $B$ under the control of the problem-setter Bob contains the problem-setting observable $\hat{B}$ (the number of the drawer in which the ball is hidden) of eigenstates and corresponding eigenvalues:
$|00\rangle_{B}$ and $00$; $|01\rangle_{B}$ and $01$; $|10\rangle_{B}$ and $10$; $|11\rangle_{B}$ and $11$.
Measuring $\hat{B}$ in a quantum superposition (or, indifferently, mixture) of all the possible problem-settings selects one of them at random.
A register $A$ under the control of the problem-solver Alice contains the number of the drawer opened by her. The content of register $A$ is the observable $\hat{A}$ of eigenstates and corresponding eigenvalues:
$|00\rangle_{A}$ and $00$; $|01\rangle_{A}$ and $01$; $|10\rangle_{A}$ and $10$; $|11\rangle_{A}$ and $11$.
Measuring $\hat{A}$ at the end of the quantum algorithm yields the number of the drawer with the ball selected by the initial measurement, namely the solution of the problem. Note that $\hat{B}$ and $\hat{A}$ commute.
The quantum description of the extended quantum algorithm is:
$\begin{array}[]{ccc}\begin{array}[]{c}\;\text{time }t_{1}\text{, meas. of}\;\hat{B}\end{array}&t_{1}\rightarrow t_{2}&\text{time }t_{2}\text{, meas. of}\;\hat{A}\\ &&\\ \begin{array}[]{c}\left(|00\rangle_{B}+|01\rangle_{B}+|10\rangle_{B}+|11\rangle_{B}\right)\\ \left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)\end{array}&&\\ \Downarrow&&\\ |01\rangle_{B}\left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)&\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow&|01\rangle_{B}|01\rangle_{A}\end{array}$ (Table I)
Here and in the following we disregard normalization.
We should start from the quantum state in the upper-left corner of Table I diagram and then follow the vertical and horizontal arrows. Initially, at time $t_{1}$, both registers, each by itself, are in a quantum superposition of all their basis vectors. By measuring $\hat{B}$ (the content of $B$) in this initial state, Bob selects the number of the drawer with the ball at random; say it comes out $01$: see the state of register $B$ under the vertical arrow [^3] By the way, Bob would be free to unitarily change at will the sorted out number of the drawer with the ball. We omit this operation that would change nothing here. .
It is now Alice’s turn. The unitary part of her problem-solving action—by the unitary transformation $\hat{\mathbf{U}}_{1,2}$—produces the solution of the problem, of course, never changing the problem-setting.
Note that we will never need to know $\hat{\mathbf{U}}_{1,2}$ (i.e. Grover algorithm). We will only need to know that, between the outcome of the initial measurement and that of the final measurement, there is a unitary evolution that builds up one-to-one correlation between the two initially unrelated observables $\hat{B}$ and $\hat{A}$.
At the end of Alice’s problem-solving action, at time $t_{2}$, register $A$ contains the solution of the problem, i.e. the number of the drawer with the ball—bottom-right corner of the diagram.
Eventually, by measuring $\hat{A}$ (the content of register $A$), which is already in one of its eigenstates, Alice acquires the solution with probability 11—the quantum state in the bottom-right corner of the diagram naturally remains unaltered throughout this measurement.
(2) Relativizing the extended quantum description of the quantum algorithm with respect to the problem-solver
We must relativize the quantum description of the quantum algorithm of Table I, which is with respect to the customary external observer, with respect to the problem-solver. This amounts to concealing from her the projection of the quantum state associated with the initial measurement of the problem-setting
Since Alice’s problem-solving action never changes the number of the drawer with the ball selected by the initial measurement and the observables $\hat{B}$ and $\hat{A}$ commute, describing this concealment is simple. It suffices to postpone after the end of Alice’s problem-solving action the projection of the quantum state associated with the initial measurement of $\hat{B}$ (or, indifferently, to postpone the very measurement of $\hat{B}$). The result is the following quantum description of the quantum algorithm with respect to Alice:
$\begin{array}[]{ccc}\begin{array}[]{c}\text{time}\text{ }t_{1}\text{, meas. of}\;\hat{B}\end{array}&t_{1}\rightarrow t_{2}&\text{\text{time }\text{}}t_{2}\text{, meas. of}\;\hat{A}\\ &&\\ \begin{array}[]{c}\left(|00\rangle_{B}+|01\rangle_{B}+|10\rangle_{B}+|11\rangle_{B}\right)\\ \left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)\end{array}&\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow&\begin{array}[]{c}|00\rangle_{B}|00\rangle_{A}+|01\rangle_{B}|01\rangle_{A}+\\ |10\rangle_{B}|10\rangle_{A}+|11\rangle_{B}|11\rangle_{A}\end{array}\\ &&\Downarrow\\ &&|01\rangle_{B}|01\rangle_{A}\end{array}$ (Table II)
To Alice, the initial measurement of $\hat{B}$ leaves the initial state of register $B$ unaltered. At the beginning of her problem-solving action she is completely ignorant of the number of the drawer with the ball selected by Bob—see the state of register $B$ on the left of $\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow$.
The unitary transformation $\hat{\mathbf{U}}_{1,2}$ is then performed for a quantum superposition of the four possible numbers of the drawer with the ball. The state at the end of it—on the right of the horizontal arrows—is a quantum superposition of tensor products, each a possible number of the drawer with the ball in register $B$ tensor product the corresponding solution (that same number) in register $A$. Since $\hat{B}$ and $\hat{A}$ commute, the final measurement of $\hat{A}$ must project this superposition on the tensor product of the number of the drawer with the ball initially selected by Bob and the corresponding solution—bottom-right corner of Table II diagram.
Note that this relativization relies on relational quantum mechanics [28, 29] where the quantum state is relative to the observer. In the following, we will call this quantum description of the quantum algorithm with respect to the problem-solver its ordinary quantum description.
(3) Time-symmetrizing the ordinary quantum description of the quantum algorithm
As anticipated in the Introduction, the ordinary quantum description of Grover algorithm is incomplete and is completed by time-symmetrizing it. We should require that the selection of the information that specifies the sorted out pair of correlated measurement outcomes, the problem-setting and the corresponding solution, evenly shares between the initial measurement of the problem setting and the final measurement of the solution. The half information selected by the initial measurement should propagate forward in time by the unitary computation of the solution $\hat{\mathbf{U}}_{1,2}$ until becoming the state immediately before the final measurement. Time-symmetrically, the complementary selection performed by the final measurement should propagate backward in time by the inverse of the computation of the solution $\hat{\mathbf{U}}_{1,2}^{{\dagger}}$ until replacing the previous outcome of the initial measurement.
We consider the time-symmetrization instance where the initial measurement of $\hat{B}$ reduces to the measurement of $\hat{B}_{l}$, the left bit of the two-bit number contained in register $B$; say that it randomly selects left bit 0. The final measurement of $\hat{A}$ should correspondingly reduce to that of $\hat{A_{r}}$, the right bit of the two-bit number contained in register $A$; say that it randomly selects right bit $1$. The randomly selected number of the drawer with the ball is thus $01$.
In this particular time-symmetrization instance, the ordinary quantum description of the quantum algorithm of Table II becomes:
$\begin{array}[]{ccc}\begin{array}[]{c}\;\text{time }t_{1}\text{, meas. of}\;\hat{B}_{l}\end{array}&t_{1\rightleftarrows}t_{2}&\text{time }t_{2}\text{,meas. of}\;\hat{A_{r}}\\ &&\\ \begin{array}[]{c}\left(|00\rangle_{B}+|01\rangle_{B}+|10\rangle_{B}+|11\rangle_{B}\right)\\ \left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)\end{array}&\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow&\begin{array}[]{c}|00\rangle_{B}|00\rangle_{A}+|01\rangle_{B}|01\rangle_{A}+\\ |10\rangle_{B}|10\rangle_{A}+|11\rangle_{B}|11\rangle_{A}\end{array}\\ &&\Downarrow\\ \begin{array}[]{c}\left(|01\rangle_{B}+|11\rangle_{B}\right)\\ \left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)\end{array}&\Leftarrow\hat{\mathbf{U}}_{1,2}^{{\dagger}}\Leftarrow&|01\rangle_{B}|01\rangle_{A}+|11\rangle_{B}|11\rangle_{A}\end{array}$ (Table III)
The projection of the initial state of register $B$ due to the measurement of $\hat{B}_{l}$—top-left corner of the diagram—is postponed after Alice’s problem-solving action, namely outside the present table which is limited to that action. With this postponement, the initial state of register $B$ goes unaltered through the initial measurement of $\hat{B}_{l}$ becoming, by $\hat{\mathbf{U}}_{1,2}$, the quantum superposition of tensor products above the vertical arrow. Alice’s final measurement of $\hat{A_{r}}$, which in this time-symmetrization instance selects the value $1$ of the right bit of register $A$, projects the state above the vertical arrow on the state below it. Then (in causal order), this final measurement outcome propagates backward in time by $\hat{\mathbf{U}}_{1,2}^{{\dagger}}$ until becoming the definitive outcome of the initial measurement. This backward in time propagation is an instance of the time-symmetrized quantum algorithm. As we can see from the diagram, the new outcome of the initial measurement tells the problem-solver, before she begins her problem-solving action, one of the possible halves of the information about the solution, namely that the ball is in drawer either 01 or 11: look at the state of register $B$ in the bottom left corner of the table.
Note that, for the mathematical equivalence between $\Leftarrow\hat{\mathbf{U}}_{1,2}^{{\dagger}}\Leftarrow$ and $\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow$, this bottom line of Table III diagram (an instance of the time-symmetrized quantum algorithm) mathematically can also be read in the usual left to right way:
$\begin{array}[]{ccc}\text{time }t_{1}&t_{1}\rightarrow t_{2}&\text{time }t_{2}\\ &&\\ \begin{array}[]{c}(|01\rangle_{B}+|11\rangle_{B})\\ \left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)\end{array}&\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow&|01\rangle_{B}|01\rangle_{A}+|11\rangle_{B}|11\rangle_{A}\end{array}$ (Table IV)
We can see that, in each instance of the time-symmetrized quantum algorithm, the computational complexity of the problem to be solved by Alice quadratically reduces from the problem of locating the ball in one of $4$ drawers to that of locating it in $2=\sqrt{4}$ drawers. In still equivalent terms, the dimension of the Hilbert space in which the search for the solution takes place quadratically reduces from $4$ basis vectors to $2=\sqrt{4}$ basis vectors.
In summary, the causal loop generated by the time-symmetrization process is the zigzag diagram of Table III. We can see that it is one of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup. The state of register $B$ in the top-left part of the diagram tells us that Alice, before beginning her problem-solving action, is completely ignorant of the number of the drawer with the ball. Through the zig-zag, this state changes into the state of register $B$ in the bottom-left part of the diagram, which tells us that Alice, before beginning her problem-solving action, knows in advance that the ball is in drawer either $01$ or $11$; we are dealing with an effective causal loop that changes its own past. She can then solve the problem in an optimal logical way by opening just one of the two drawers, indeed as in the classical logic description of what the quantum algorithms does.
Also note that table III diagram goes along with the fact that the unitary part of the quantum algorithm, $\hat{\mathbf{U}}_{1,2}$, takes the opening just one drawer, but does not imply it. We know this fact from outside Table III diagram: we know that this is what the unitary part of Grover algorithm takes. The diagram says that the problem-solver knows in advance that the ball is in a pair of drawers rather than four—bottom left corner of the diagram—and this could allow her to solve the problem by opening a single drawer—bottom line of the diagram of table III or the diagram of table IV. The advanced knowledge rule tells us the room for the quantum speedup but does not imply it.
Eventually, we should take the quantum superposition of all the time-symmetrization instances.
Note that this time-symmetrization, mathematically, leaves the unitary computation of the solution, $\hat{\mathbf{U}}_{1,2}$, unaltered, but changes the behavior of causality along it. The single causal loop of Table II is replaced by the quantum superposition of the causal loops whose mutually exclusive or is logically implied by the sheer existence of the quantum speedup.
In subsection 2.6, we have seen that all this, first, immediately extends to any number of drawers, then, to more general forms of quantum correlation.
Acknowledgements
Thanks are due to Eliahu Cohen, Artur Ekert, Avshalom Elitzur and David Finkelstein for useful comments, Daniel Sheehan for organizing the San Diego AAAS Pacific Division conferences on retrocausality, a far-looking forum for the discussion of frontier, also unorthodox, physics, and Mario Rasetti for organizing the first series of international workshops on quantum communication and computation, the 1991–96 Elsag Bailey–ISI Turin Villa Gualino workshops.