### v7

# Quantum mechanics provides the physical basis of teleological evolutions

Abstract
We show that the quantum computational speedup of quantum algorithms is due to their teleological character, their being evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again). We also show that, under the quantum cosmological assumption and for the Fine-tuned Universe version of the Anthropic Principle, the physical basis of the teleological character of quantum algorithms applies as well to the evolutions of the living for which the teleological notion was originally conceived.
1 Introduction
Once applied to the evolutions of the living by natural philosophy, the notion of teleological evolution—of an evolution attracted by a final goal—was dismissed with the advent of modern science for the alleged lack of a physical basis. For its finalistic character, it was also ascribed to the category of beliefs. In this work, we shall show that the relatively recent discovery of the quantum computational speedup, the fact that quantum algorithms can solve problems with fewer logical operations (computation steps) than the minimum classically possible, should change this position of science. Indeed, we shall show that the quantum speedup of quantum algorithms is due to their teleological character.
In previous works of coauthors and ours [1, 2, 3, 4, 5, 6], we developed an unconventional, time-symmetric approach to the quantum speedup. In particular, in [4], we showed that:
In quantum algorithms, it is as if the problem-solver, before beginning her problem-solving action, knew in advance any of the possible halves of the information that specifies the solution of the problem she will produce and measure in the future and could use this knowledge to compute the solution with fewer computation steps; the number of computation steps needed to solve the problem in an optimal quantum way is that needed to solve it in an optimal logical way benefiting from the advanced knowledge in question.
We have called the above the advanced knowledge rule. In the above publications, it was given as an empirical rule holding for all the quantum algorithms examined, which comprised the major ones. In the present work, we should have demonstrated it. Note that this rule, in addition to providing a unified quantitative explanation of all quantum speedups that is absent in mainstream quantum information, implies the teleological character of quantum algorithms. Indeed, according to it, they are evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce and measure in the future (the solution of the problem again): Without parentheses and their contents, this can well be considered the definition of teleological evolution. This also clarifies the finalistic character of the notion. A teleological evolution is not attracted by some superior scientifically undefinable goal, as argued by the opponents of the notion, but, quite simply, by its own end product.
Note that the advanced knowledge rule establishes the number of computation steps needed to solve an (oracle) problem in an optimal quantum way. Whether the quantum algorithm is optimal is another thing. We can say that the advanced knowledge rule tells us the room for the quantum speedup.
Of course, the advanced knowledge rule, its necessarily implying a form of retrocausality, is a destabilizing notion. Let us address in the first place this critical aspect of the work.
Knowing in advance, on the part of the problem-solver, any of the possible halves of the information about the solution she will produce and measure in the future and using this knowledge to produce the solution with fewer computation steps is obviously a causal loop. As there is a plurality of ways of taking half of the information about the solution, we can say that, according to the advanced knowledge rule, the quantum speedup is explained by the mutually exclusive or of these causal loops.
As far as we know, causal loops are completely absent from ordinary non-relativistic quantum mechanics. Of course, the current quantum description of quantum algorithms is causal in character—i.e. with causality always going in the same time-direction. Naturally, also this quantum description explains the quantum speedup of each quantum algorithm, although in a different way for each quantum algorithm—we shall go back to this further below.
We have to justify the discrepancy between what is described by the advanced knowledge rule and the current quantum description of quantum algorithms. It turns out that the current quantum description is incomplete and, once completed, it comprises the advanced knowledge rule. It is incomplete in a double way; the first is trivial (but should be amended first), the second is fundamental.
The trivial incompleteness consists in the fact that the quantum description is limited to the unitary computation of the solution where the quantum speedup resides and the final measurement of the latter. Of course, the canonically complete quantum description of a quantum process consists of initial measurement, unitary evolution, and final measurement. This incompleteness is also trivially amended. To the quantum description of the process of solving the problem, we should add that of the upstream process of setting the problem. An initial measurement of the problem-setting in a quantum superposition (or, indifferently, mixture) of all the possible problem-settings selects one of them at random; the unitary computation and final measurement of the corresponding solution follow [^1] We exemplify what we mean by problem-setting. Consider the following problem: Bob, the problem-setter, hides a ball in a chest of drawers, Alice, the problem-solver, should locate it by opening drawers. The number of the drawer with the ball selected by Bob is the problem-setting.. We call this the ordinary quantum description of the quantum algorithm; “ordinary” because it is still the causal quantum description of mainstream science.
This ordinary quantum description is still incomplete, this time in the fundamental way. Further below, we shall see that it is exactly the same incompleteness lamented by Einstein et al in the case of the quantum description of quantum nonlocality [7].
Before proceeding, let us simplify the situation by requiring that the solution of the problem is an invertible function of the problem-setting and is always determined with certainty (of course, the problem-setting and the solution are two commuting observables). We should also keep in mind that the unitary computation of the solution never changes the problem-setting.
With the above, we can go to the amendment of the fundamental incompleteness of the quantum description we are dealing with. It turns out that this is done by time-symmetrizing it. It is a time-symmetrization inspired to time-symmetric quantum mechanics [8, 9, 10]. As we shall see, the time-symmetrization in question also turns out to be mandatory to satisfy an atemporal symmetry implicit in quantum algorithms and highlighted by their present (ordinary) quantum description.
Things go as follows. Ordinarily, the information that specifies the problem-setting and the corresponding solution is all selected by the initial measurement of the problem-setting. We should instead require that this selection evenly shares between the initial measurement of the problem-setting and the final measurement of the solution. As there is a plurality of ways of evenly sharing, we should take their quantum superposition.
In each time-symmetrization instance, the outcome of the initial measurement should propagate forward in time by the unitary computation of the solution until it becomes the state immediately before the final measurement. Time-symmetrically, the outcome of the final measurement should propagate backward in time, by the inverse of the unitary computation of the solution, until replacing the previous outcome of the initial measurement. As we shall see, this backward in time propagation tells the problem solver, before she begins her problem-solving action, one of the possible halves of the information about the solution. Indeed, the selection performed by the final measurement affects the initial measurement outcome, changing the initial complete ignorance of the solution on the part of the problem-solver into knowledge of one of the possible halves of the information about it [^2] On the possibility that a future selection affects a past measurement outcome see [10].. She can then use this knowledge to produce the solution with fewer computation steps. In other words, this time-symmetrization generates the quantum superposition of the causal loops whose mutually exclusive or is implicit in the advanced knowledge rule. Besides reconciling the ordinary quantum description of the quantum algorithms with the advanced knowledge rule, this is the demonstration of the latter—we should keep in mind that the time-symmetrization we are dealing with is mandatory.
By the way, we shall provide another, independent, demonstration of the advanced knowledge rule. The fact that the ordinary quantum description of the quantum algorithm involves only the measurements of commuting observables (the problem-setting and the solution of the problem) makes it subject to classical logic. Now, as we shall see, the classical logic description of the fact that the quantum algorithm computes the solution with fewer computation steps than the minimum possible in the classical case very directly implies the advanced knowledge rule. It is interesting to provide also this demonstration because it very clearly highlights the interplay between time-forward and time-backward causality in unitary quantum evolutions that produces the quantum speedup.
Let us make another specification in view of what will follow. Of course, the sending of information backward in time operated by the time-symmetrization of the ordinary quantum description of the quantum algorithm must be along a time-reversible process. Here, it is the unitary computation of the solution. The special character of its time-reversibility should be underlined. Unlike the classical time-revesibility, which is necessarily an idealization, the present kind is both ideal and objective. This is, of course, an essentially quantum feature that stands at the basis of the present time-symmetric explanation of the quantum speedup.
We have seen that demonstrating the advanced knowledge rule also amounts to demonstrating the teleological character of quantum algorithms. Of course, this would already falsify the current opinion that teleological evolutions cannot have a physical basis. However, their present physical basis is necessarily in the quantum world and the teleological notion was applied to the evolutions of the living. We shall show that, under the quantum cosmological assumption and for the Fine-tuned Version of the Anthropic Principle, the physical basis of the teleological character of quantum algorithms identically applies to the evolutions of the living. Under the quantum cosmological assumption, the evolution of the universe toward life becomes quantum. For that version of the Anthropic principle, it acquires the form of a quantum algorithm and, as such, can be endowed with quantum speed up and consequently be teleological in character. The evolution of life on Earth, in principle the trace over all the rest of the former evolution, and, with it, the evolutions of life, would inherit that possible teleological character.
In summary, the time-symmetrization we are dealing with (and the corresponding completion of the quantum description) turns out to have a special unification power. Besides providing a unified quantitative explanation of all quantum speedups (the one provided by the advanced knowledge rule), it unifies the explanation of the quantum speedups with that of quantum nonlocality and, under the quantum cosmological assumption, it unifies the teleological character of quantum algorithms with that of the (macroscopic) evolutions of the living for which the teleological notion was originally conceived.
2 Time-symmetric explanation of the quantum speedup
We provide a description of quantum algorithms and their quantum speedups open to the time-symmetric coexistence of causality and retrocausality along unitary quantum evolutions. In this section, we provide it at a conceptual level, but making reference to the formalizations postponed in an Appendix. The section is divided into subsections corresponding to the various arguments.
2.1 Critique of the current way of viewing quantum computation
Let us first acknowledge that that of the quantum computational speedup has been one of the fundamental discoveries of the past century. Together with that of quantum communication, it has created an entirely new branch of physics. However, for the sake of clarity, we will be completely frank in criticizing what we think are the shortcomings of the current way of viewing quantum computation, likely related to the rejection of the notion of retrocausality.
We repeat here for convenience the advanced knowledge rule; of course, it plays an essential role in the present work.
In quantum algorithms, it is as if the problem-solver always knew in advance, before beginning her problem-solving action, any of the possible halves of the information that specifies the solution of the problem she will produce and measure in the future and could use this knowledge to produce the solution with fewer computation steps. The number of computation steps needed in the optimal quantum case is that needed in the optimal logical case benefiting from the advanced knowledge in question.
This rule would offer a unified quantitative explanation of the quantum speedups that is still lacking in mainstream quantum information. Let us discuss this point. To the best of our knowledge, this absence has never been acknowledged by the scientific community.
People currently say that there is an explanation of the quantum speedup; it would be that of quantum parallel computation, the fact that many computations are performed at the same time in quantum superposition. Indeed, this captures the essential character of quantum computation, but, being a qualitative explanation of the quantum speedup, which is a quantitative feature, it cannot be considered its explanation. Of course, science essentially differs from opinion because of its quantitative character; then, settling for a qualitative explanation of a quantitative feature is rather unusual. This is our opinion, of course. However, as a matter of fact, outside the present retrocasusal approach to the quantum speedup, in the literature, there is no trace whatsoever of any search for a unifying-quantitative explanation of it.
Neither can be considered a satisfactory explanation of the quantum speedup the fact that each quantum algorithm naturally explains its quantum speedup. In fact, it would be an explanation completely different for different quantum algorithms. It would be like having different explanations for different throws of a stone without knowing that they all describe a parabola. Naturally, unification of explanations is a fundamental objective of science.
Of course, the advanced knowledge rule would exactly be that unifying-quantitative explanation. In this section, we shall provide the two demonstrations of it, beginning with the one relying on the classical logic description of the quantum speedup.
2.2 First demonstration of the advanced knowledge rule
Let us recall the quantum description of quantum algorithms extended to the process of solving the problem already provided in the Introduction:
An initial measurement in a quantum superposition (or, indifferently, mixture) of all the possible problem-settings selects one of them at random [^3] The problem-setter could unitarily change it at will; this, here, would change nothing.. The unitary computation and the final measurement of the corresponding solution follow.
For the sake of simplicity, we consider the case that the solution is an invertible function of the problem-setting and is determined with certainty. Table I of Section (1) of the Appendix provides the formalization of this extended quantum description in the simplest instance of Grover algorithm. Here, the problem is as follows: Bob, the problem-setter, hides a ball in a chest of four drawers, Alice, the problem-solver, is to locate it (i.e. identify the number of the drawer with the ball) by opening drawers.
Note that the quantum description of the quantum algorithm of Table I is with respect to the customary observer external to the quantum process. To present ends, we should relativize it with respect to the problem-solver who, for the fact of computing the solution of the problem, is part of the process. Of course, the advanced knowledge of one of the possible halves of the information about the solution we are dealing with will have to appear in the (completed) quantum description of the quantum algorithm with respect to her. Now, the problem-setting selected by the initial measurement should be concealed from her: as we shall see, it would tell her the solution of the problem before she begins her problem-solving action: Of course, at this time, instead, she is completely ignorant of the solution. Since the unitary computation of the solution never changes the problem-setting selected by the initial measurement, this concealment is easy. It suffices to postpone the projection of the quantum state associated with the initial measurement to the end of her unitary computation of the solution; see Table II of Section (2) of the Appendix.
Let us call this extended quantum description with respect to the problem-solver “the ordinary quantum description of the quantum algorithm”. It is ordinary because it is still the customary “causal” quantum description of mainstream quantum mechanics—where causality always goes in the same time-direction.
We can now go to the first demonstration of the advanced knowledge rule. It simply consists in describing in a classical logic way what the quantum algorithm does—computing the solution of the problem with a number of computation steps that are less than the minimum possible in the classical case.
By the way, since their initial and final measurements are of commuting observables (the problem-setting and the solution of the problem) quantum algorithms, and what they do, should be subject to classical logic.
Now, we have to choose the quantum algorithm to work with. We need a quantum algorithm that certainly gives a quantum speedup with respect to what is classically possible. Now, most quantum algorithms yield a quantum speedup with respect to their best classical counterpart, but whether the latter is the best possible is not known. Excluding elementary one-computation-step quantum algorithms whose dimensions do not scale up, we find our choice limited to Grover’s quantum search algorithm [13]. Besides yielding a well determined quantum speedup with respect to what is classically possible, it has the following convenient features: (i) In Long’s version [14], it always provides the solution with certainty; this simplifies things. (ii) In it, the solution is an invertible function of the problem-setting; this simplifies its time-symmetrization. (iii) It is optimal in character; this, of course, constitutes an important reference. (iv) It is a most universal quantum algorithm: it can easily be adapted to solve any NP problem with a quadratic quantum speedup; more than a single quantum algorithm, it is an entire test bed on which to test the present theory of the quantum speedup.
Let us show that (under classical logic) the quantum speedup over what is classically possible delivered by Grover algorithm logically implies the advanced knowledge rule. We begin with the simplest instance of this quantum algorithm: everything immediately generalizes first to the full fledged Grover algorithm and then to most general forms of quantum correlation (in quantum algorithms, that between the problem-setting and the solution of the problem).
The problem to be solved is as follows. Bob, the problem-setter, hides a ball in a chest of four drawers. Alice, the problem-solver, is to locate the ball by opening drawers. In the worst classical case, she has to open three drawers: If the ball is only in the third drawer opened, she has solved the problem. If not, it must be in the only drawer not yet opened, and she has solved the problem of locating the ball as well.
Instead, Grover algorithm (its simplest instance) always solves the problem by opening just one drawer, in a quantum superposition of the four possible drawer numbers. Obviously, always solving the problem by opening just one drawer would be impossible in the classical case: indeed, we are dealing with a quantum computational speedup.
There is a classical logic equivalent of the fact that the quantum algorithm always solves the problem by opening a single drawer. By the way, since the fact that a single drawer is opened is logically true in itself, we can forget the specification that this is done for a quantum superposition of the four drawer numbers.
The logical equivalence is as follows. Logically, it is as if the problem-solver always knew in advance, before beginning her problem-solving action, that the ball is in a given pair of drawers (in a mutually exclusive way one of the three pairs of drawers in which the ball must be) and took advantage of this knowledge to solve the problem by opening just one of them. In fact, if the ball is in the opened drawer, she has solved the problem; if not, it must be in the other drawer and she has solved the problem of locating the ball as well, in either case having opened just one drawer. By the way, this also means knowing in advance half of the information that specifies the solution of the problem that will be produced and measured in the future.
Let us open a parentheses to observe that we are not talking about the quantum algorithm as it were a classical algorithm. In other words, here, classical logic should not be confused with classical physics, which might be a temptation in the present case.
Always knowing in advance that the ball is in a given pair of drawers is logically necessary to always solve the problem by opening just one drawer. Then the question becomes: from where does Alice, before beginning her problem-solving action, always get that information about the solution? Since, for her, the only possible source of it is the solution she will produce and measure in the future, the only possible answer is that the information in question retrocausally comes to her from her future measurement of the solution. This is formalized in Section (3) of the Appendix. However, before looking at it, one should read the description of the time-symmetrization of the quantum algorithm that we shall provide in a moment.
Note that we are dealing with a causal loop. Alice knows in advance one of the possible halves of the information about the solution she will produce and measure in the future (that the ball is in two rather than four drawers) and can use this knowledge to produce the solution by opening just one drawer. There is a causal loop for each way of taking half of the information about the solution. Therefore, we can say that this quantum speedup is logically explained by the mutually exclusive or of the causal loops we are dealing with. Of course, this is the advanced knowledge rule for the present instance of Grover algorithm, logically implied by the sheer existence of the quantum speedup.
This explanation of the quantum speedup immediately generalizes to any number $N$ of drawers. As is well known, the number of drawer openings needed by the (optimal) Grover algorithm is on the order of $\sqrt{N}$ against the order of $N$ of the best classical case, hence its quadratic quantum speedup. This is logically equivalent to knowing in advance, in a mutually exclusive way, that the ball is in one of the possible tuples of $\sqrt{N}$ drawers and benefiting from this knowledge to locate the ball in the best logical way by opening the order of $\sqrt{N}$ drawers. Of course, this is what the optimal Grover algorithm does. Also in this case, using the advanced knowledge of half of the information about the solution in an optimal logical way yields the optimal quantum speedup of Grover algorithm. This is, of course, the advanced knowledge rule for the full fledged Grover algorithm.
Note that, in equivalent terms, the computational complexity of the problem to be solved reduces from searching the ball in $N$ drawers to searching it in $\sqrt{N}$ drawers. In still equivalent terms, the dimension of the Hilbert space in which the search for the solution takes place reduces from $N$ to $\sqrt{N}$.
Holding for Grover algorithm, the above retrocausal explanation of the quantum speedup must hold for the process of solving with a quadratic quantum speedup any NP problem. We shall see that there are further generalizations in the following of this section. In view of them, note that, in order to derive the present explanation of the quantum speedup, we never needed to know how Grover algorithm achieves the latter, we only needed to know that it achieves it.
By the way, let us underline the fact that the advanced knowledge rule tells us what can be the quantum speedup in the optimal quantum case. Of course, whether this quantum speedup is achieved, or achieved in part, or not at all, depends on the character of the unitary evolution that computes the solution. Therefore, we can say that the advanced knowledge rule tells us the room for the quantum speedup. Incidentally, achieving a quantum speedup may not be difficult: [15] shows that even a random quantum search for the solution yields just a suboptimal quantum speedup.
2.3 Completing the ordinary quantum description of quantum algorithms by time-symmetrizing it
The above subsection also says that the ordinary quantum description of quantum algorithms, which obviously does not describe the advanced knowledge rule, is incomplete. Of course, to bring it in line with their classical logic description, it should be completed. It turns out that this is simply done by time-symmetrizing it; we shall also see that this time-symmetrization is mandatory to satisfy an atemporal symmetry implicit in this quantum description.
We describe this operation. For the sake of simplicity, we consider the case that the solution is determined with certainty and is an invertible function of the problem-setting—exactly as in Grover algorithm.
Naturally, in the ordinary quantum description of the quantum algorithm, the initial measurement of the problem-setting selects the information that specifies both the problem-setting and the corresponding solution. To time-symmetrize this ordinary quantum description, we should evenly share the selection of this information between the initial measurement of the problem-setting and the final measurement of the solution. Since there is a plurality of possible ways of evenly sharing, we should take their quantum superposition.
In each time-symmetrization instance, the selection performed by the initial measurement propagates forward in time by the unitary computation of the solution until it becomes the state immediately before the final measurement. Time-symmetrically, the selection performed by the final measurement propagates backward in time by the inverse of the unitary computation of the solution until replacing the previous outcome of the initial measurement. This backward in time propagation, which inherits both selections, is an instance of the time-symmetrized quantum description. As we can see in Table III of Section (3) of the Appendix, in the initial measurement outcome of the time-symmetrized instance of the quantum description, it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future (left of the bottom line of Table III diagram); she can then use this knowledge to compute the solution with fewer computation steps (bottom line of Table III diagram or Table IV diagram). The full Table III diagram is the physical correspondent of one of the causal loops whose mutually exclusive or is logically implied by the sheer existence of the quantum speedup.
Eventually, we should take the quantum superposition of all the time-symmetrization instances. The time-symmetrized quantum description of the quantum algorithm (i.e. its completed quantum description) becomes the quantum superposition of the causal loops whose mutually exclusive or is logically implied by the quantum speedup: Thus completed, also the quantum description of the quantum algorithm describes the advanced knowledge rule described by its classical logic description—more precisely it describes its physical counterpart.
2.4 Second demonstration of the advanced knowledge rule
The second demonstration of the advanced knowledge rule simply comes from the observation that the time-symmetrization of the ordinary quantum description of the quantum algorithm, which provides the (physical counterpart of the) advanced knowledge rule, is mandatory also for another reason, as follows.
In the ordinary quantum description of the quantum algorithm, of course, the unitary computation of the solution never changes the corresponding problem-setting. Therefore, one should be free to postpone the initial measurement of the problem-setting along the subsequent unitary computation of the solution until it becomes simultaneous with the final measurement of the solution. At this point, we find ourselves with the simultaneous measurements of two maximally entangled observables (the problem-setting and the solution, one an invertible function of the other), and the time-symmetrization we are dealing with must already be in place to represent the perfect (atemporal) symmetry between these two measurements. Of course, there would be no reason why a measurement selected more information than the other. As we have just seen, this time symmetrization generates the quantum superposition of the causal loops whose mutually exclusive or is logically implied by the sheer existence of the quantum speedup—the physical counterpart of the advanced knowledge rule. This is indeed the second demonstration of the advanced knowledge rule.
By the way, this result is reminiscent of Wheeler and Feynman’s [16] time-symmetric absorber theory; we would say: with the advantage that the time-symmetrization we are presently dealing with relies on the reversibility of unitary quantum evolutions, which, unlike classical reversibility, is such in both an ideal and objective way.
2.5 Origin of the mutual causality
A quantum algorithm is a relatively complex process. A natural question is whether the mutual causality implicit in the classical logic description of its overall result, i.e. the quantum speedup, derives from some more elementary part of it. Of course, in its ordinary quantum description, the unitary computation of the solution is an alternation of generations of quantum state superpositions and quantum interferences. Since the latter have a classical analog, the essentially quantum generation of a quantum state superposition is the suspect.
Let us show that the classical logic description of the unitary generation of a quantum state superposition necessarily involves retrocausality.
Think of a photon going in quantum superposition through the two branches of a Mach-Zehnder interferometer. The two branches should be tuned in such a way that the photon always ends up in the same detector, so that we have a (of course, time-reversible) unitary evolution that goes from before the photon enters the interferometer to immediately before its detection.
The classical logic description of this quantum superposition can be as follows. In entering the interferometer, the photon takes, say, the upper branch until it reaches the state immediately before detection (by the detector in question). Of course, this also provides the phase of the photon along the path followed. Then (in causal order), the photon goes backward in time undoing the path just followed until before entering the interferometer. Then, it goes forward in time again taking the lower branch until reaching the detector in question, and this time, be detected. The zigzag path followed in this way by the photon through the interferometer, with the associated photon phases, of course, describes in a classical logic way the fact that the photon goes through the interferometer in a quantum superposition of the upper and lower branches, eventually ending up in that detector. It should be clear that the classical logic description of this unitary generation of a quantum state superposition necessarily involves the coexistence of the two time directions of causality along the same unitary process.
We should think that the classical logic description of the sequence of generations of quantum state superpositions and interferences that constitutes the unitary computation of the solution, in the end, simplifies into a description that implies the advanced knowledge rule.
Summarizing, at the root of the form of mutual-causality we are dealing with there would be the fundamental notion of quantum state superposition. By the way, note that this is not an idle speculation as it might even seem. Indeed, we are dealing with a form of coexistence of the two time-directions of causality with huge consequences, such as providing a physical basis for the notion of teleological evolution.
2.6 Generalization
In the second demonstration of the advanced knowledge rule, we have seen that the ordinary quantum description of a quantum algorithm whose solution is an invertible function of the problem-setting and is determined with certainty is incomplete and must be completed by time symmetrizing it. One can see that the necessity of this time-symmetrization does not depend on the fact that we are dealing with a quantum algorithm; there is this immediate generalization:
The ordinary quantum description of any unitary evolution between two one-to-one correlated measurement outcomes that does not change the eigenvalue selected by the initial measurement is incomplete and is completed by time-symmetrizing it.
Indeed, as one can see, to this kind of unitary evolution one can apply the argument for which it must be time-symmetrized to satisfy the atemporal symmetry implicit in it. Let us call this rule the quantum correlation time-symmetrization rule.
In the case that the unitary evolution builds up one-to-one correlation between two initially unrelated observables, as in quantum search algorithms, this building up can always occur with quadratic quantum speedup, in the sense that there is always the room for its occurring with quantum speedup. This is what the time-symmetrization of its ordinary quantum description says—see Section (3) of the Appendix. Of course, in the case where we are not dealing with a quantum algorithm, the notion of advanced knowledge of one of the possible halves of the information about the solution must be replaced by the equivalent notion of quadratic reduction of the dimension of the Hilbert space in which the search for that one-to-one correlation takes place, from $N$ to $\sqrt{N}$.
The case in which the one-to-one quantum correlation is not built up by the unitary evolution because it preexists it (more precisely, what should preexist is the maximum entanglement between the observables of the two measurements) is, for example, that of quantum nonlocality with the two measurement apparatuses aligned. Here, the unitary evolution spatially separates the subsystems of two already maximally entangled observables. With the two measurement apparatuses aligned and the spatial separation process preceding the first measurement, this responds to the quantum correlation time-symmetrization rule.
As shown in [5], the time-symmetrization of the ordinary quantum description of this quantum nonlocality yields a quantum superposition of causal loops that locally explain it and play the role of the hidden variables envisaged by Einstein et al [7]. Indeed, these causal loops locally go from one to the other measurement outcome via the time the two subsystems were not yet space separate, where they locally tell the subsystem of the not yet measured observable the state to assume in view of the future measurement. Note that this is exactly what Einstein et al required from their hidden variables. That was, of course, without resorting to any mutual-causality, but under a deterministic view of the measurement process equivalent in this case. Of course, this should also mean that the incompleteness we are presently ascribing to the ordinary quantum description of quantum algorithms is the same that Einstein and others ascribed to the quantum description of quantum nonlocality.
By the way, the fact that the present explanation of the quantum speedup also demonstrates the incompleteness of the quantum description highlighted by Einstein et al might also be destabilizing. Indeed, after highlighting for the first time quantum nonlocality and inspiring fundamental works like Bell’s theorem [17] and Ekert’s quantum cryptography based on it [18], their argument had been dismissed by the physicists’ community because of the alleged absence of a satisfactory way of completing the quantum description by means of local hidden variables. In particular, the argument of the incompleteness of the quantum description has been forgotten.
2.7 The possible teleological character of the evolutions of the living
We have seen that the possibility of a teleological character of a quantum algorithm, of its occurring with quantum speedup and thus be teleological in character, depends only on the form of its ordinary quantum description: of the unitary evolution between the initial an the final measurement we need only to know the input and the output. We do not need to know its inside. All this defines the room for the quantum speedup, which is what is of interest here. The unitary evolution might even not be a quantum algorithm.
Let us show how the teleological character of quantum algorithms could be ported to the (macroscopic) evolutions of the living for which the teleological notion was originally conceived. The first thing one should observe is that quantum algorithms belong to the microscopic world of quantum phenomena and, as things are now, have durations of a few dozen of microseconds: more than that, there is quantum decoherence. The evolutions of life, instead, can have the size of the earth and have duration of billions of years. The distance between the two is immense but, it should be noted, it would be so under the customary classical vision of macroscopic reality. Instead, under the quantum cosmological assumption that the universe evolves in a unitary quantum way, this distance vanishes. By the way, recall the perfect time-reversibility of unitary quantum evolutions.
Now there is a surprising fact. Under the quantum cosmological assumption and for the Fine-tuned Universe version of the Anthropic Principle, the evolution of the universe toward the value of life acquires the form of a quantum search algorithm and, as such, it can occur with quantum speedup and thus be teleological in character, this in the sense that there is certainly the room for that.
Let us explain. That version of the principle is about the extraordinary theoretical fact that the least change of the value (for values) that the fundamental physical constants has in the universe we are in would have produced a relatively trivial universe unable to develop life. Under the quantum cosmological assumption, this can be represented as follows. We should introduce two commuting observables of the same size: the value of the fundamental constants and the “value of life” ; the latter should assume, in particular, the values “absence of life” and “existence of sentient life”.
We should then consider a quantum superposition (or, indifferently, mixture) of all the possible universes, each with its own value of the fundamental constants, all obeying the same fundamental laws (but for the specific value of the fundamental constants). Of course, at the beginning of time, in this quantum superposition, these two observables must be unrelated: no matter the value of the fundamental constants, the value of life is “absence of life”. We should now consider the unitary evolution of this quantum superposition of all possible universes toward the value of life. At the time of today, this evolution must have built up maximal entanglement (say, one-to-one correlation) between the value of the fundamental constants since the beginning of time and the value of life, of course, without ever changing the former value. By the way, according to cosmology, at the time of today, the absence of life in the universes that cannot produce it would already be definitive.
We have seen that this kind of building up of quantum correlation can always occur with quantum speedup (in the sense that there is always the room for it) and therefore be teleological in character. In the universe we are in, the evolution of the universe toward the value of life becomes evolution of the universe toward life; and, at the time of today, there are the sentient observers who measure the value of life, also selecting the value of the fundamental constants maximally entangled with it. By the way, since the latter value is never changed by the unitary evolution of the universe toward life, all is as if it was measured at the beginning of time. We could also think that, if the search for the value of life by the evolution of the universe is random, it would occur with quantum speedup.
Summarizing, the evolution of the universe toward life could occur with quantum speedup and consequently be teleological in character. Let us show that this possibility would then be inherited by the evolution of life on Earth and the evolution of species.
In principle, the evolution of life on Earth can be derived from the evolution of the universe toward life by tracing out all the rest. Conversely, we can speak of an image of the evolution of life on Earth in the evolution of the universe toward life. Then, first, evolution times must be the same for the evolution of life on Earth and its image in the evolution of the universe. Second, the amount of physical resources used by the evolution of life on Earth cannot exceed the amount of physical resources used by its image in the evolution of the universe. In fact, the latter resources comprise the former.
In summary, under the quantum cosmological assumption, the evolution of life on Earth could occur with quantum speedup (in the sense that there would always be the room for it) and consequently be teleological in character; it might actually occur with quantum speedup if the search for life by the evolution of the universe toward life were random. Then, just for consistency, the evolutions of species would also be the same.
Not to hide behind a finger, let us exemplify what a teleological character of the evolution of life would imply; of course, it would be very destabilizing with respect to the present way of viewing both the evolutions of the living and physical reality.
It is known that, for many millions of years, the ancestors of the birds were feathered dinosaurs that did not fly. Very light, strong but flexible, cave and air-proof, from the fossils found, even at that time feathers were ideal for flight, very close to those of the future birds. So, why anticipate something designed for flight many millions of years in advance?
The usual explanation, of course, is causal. First, an ancestor of feathers appeared by random DNA mutation, then evolved as a good thermal insulator and perhaps to speed escape from predators; only eventually (by sheer coincidence) feathers turned out to be what was needed to conquer the air.
All this (without the parentheses and its content) could remain true in the teleological view with just one, but very significant, addition: Those random mutations that would have led in the future to the conquest of the air would have more likely occurred. In our possibly biased opinion, this mutual-causality explanation of the anticipation in question would be much more convincing than the customary causal explanation. Indeed, the feathers of the fossils of those non-flying dinosaurs are almost identical to those of their flying descendants. Our preprint [6] provides a more extensive discussion of the implications of the teleological notion.
Another consequence of the present theory is that, according to it, a teleological evolution is an evolution endowed with quantum speedup. Of course, this is a feature of teleological evolutions completely absent in their historical notion. It could contribute to explaining why some steps of the evolutions of the living have been so rapid to defy Darwinian gradualism, see [19]. As is well known, have been of this kind, in particular, the Cambrian Explosion and modern ecological adaptation.
Of course, being under the quantum cosmological assumption, all the above is hypothetical. This is to decouple it from the teleological character of quantum algorithms, which we would consider watertight.
3 Discussion
First, let us show that the abandonment of the teleological perspective in the natural sciences was and is not without criticism.
An example is a famous sentence by the evolutionary biologist Haldane [23]:
Teleology is like mistress to a biologist, he cannot live without her, but is unwilling to be seen with her in public.
The other is the position of the natural philosopher Nagel [24], who provides a series of very practical arguments for which the mechanistic, neo-Darwinian conception of nature could not account for the fundamental features of the living, and this in favor of a holistic and teleological view.
It could be said that, in general, natural philosophy is still pro the teleological notion. The following is a quotation from the natural philosophers Allen and Neal [22]:
Teleological notions are largely considered ineliminable from modern biological sciences, such as evolutionary biology, genetics, medicine, ethology, and psychiatry, because they play an important explanatory role.
Against the above, there is the position of the physical science and its de facto rejection of the teleological notion. By the way, this is also the position officially adopted by the natural sciences.
In our possibly biased view, natural philosophy would respond better to empirical evidence and be less affected by what we would consider a de facto prejudice of science.
Note that a teleological character of the evolutions of the living on the Earth could be further explored by investigating, either theoretically or by means of toy models, the character that the operation of “tracing out all the rest” would acquire in the present context. We should consider the unitary quantum evolution of a system endowed with quantum speedup, standing for the evolution of the universe toward life, and derive from it, by tracing out all the rest, the almost autonomous evolution of a subsystem of it, standing for the evolution of life on Earth. We should look for a way of obtaining an evolution of the subsystem that is teleological (thus quantum) in the longer term and (about) classical in the shorter term. If this were possible, it would go in favor of the quantum cosmological reading of the Anthropic Principle that we have proposed. By the way, this could also be another perspective under which to address the problem of unifying quantum and classical theory, in this case under the present form of time-symmetric quantum mechanics.
The same might also hold for cosmology. After all, under the quantum cosmological assumption, also the physical part of the evolution of the universe toward life should be subject to the Fine-tuned Universe version of the Anthropic Principle. Also this part could be endowed with quantum speedup and thus be teleological in character. It might be interesting to confront this possibility with the present cosmological notions, naturally comprising their open problems.
We have already noted that the present notion of teleological evolution implies the quantum speedup of the latter. This would, of course, be a novelty with respect to the historical notion of teleological evolution. It might even better explain than Darwinian gradualism the extraordinary speed with which some evolution steps occurred.
Another question is how to describe a macroscopic evolution of life endowed with quantum speedup. Of course, classical physics would turn out to be unable to describe not only the processes of the microscopic world but also the teleological evolutions of the macroscopic world.
We quote a precedent in the quantum description of macroscopic natural evolutions. It is about the asymptotic behavior of predator species competing for the same prey [25]. It is explained by a quantum model, whereas it would be difficult to explain it in a classical way. In our case, the canonical way of describing the evolution of life on Earth would be to obtain it from the description of the evolution of the universe toward life by tracing-out all the rest. Since this obviously holds only in principle, we should also model the macroscopic teleological evolutions of species endowed with quantum speedup in a quantum way. In other words, we should accept quantum superposition, entanglement, and correlation between the observables of macroscopic objects.
This might even be a good approximation, since, in the universe we are in, the evolution of life on Earth in its time lapse should have been almost autonomous from the rest of the unitary quantum evolution of the universe toward life. Indeed, if it were completely autonomous, of course, it would be a unitary quantum evolution itself.
4 Conclusions
The present work goes against a long-standing de facto no retrocausality tradition of physics. More specifically, since causality could well go backward in time along a unitary quantum evolution that is both objectively and ideally time-reversible, we should say that, in the quantum world, the scientific tradition amounts to asserting the mutual exclusivity between the two possible time directions of causality.
Let us first address the reliability of the present unconventional work.
The logic for which quantum algorithms should be teleological evolutions is simple. Always locating a ball hidden in one of four drawers by opening a single drawer, as the quantum algorithm does, is obviously equivalent to knowing in advance that the ball is, in a mutually exclusive way, in one of the three pairs of drawers in which it must be. This, of course, implies knowing in advance, in a mutually exclusive way, one of the possible halves of the information about the solution one will produce and measure in the future and using this knowledge to produce the solution with fewer computation steps. This attraction from the future is indeed the teleological character of the quantum algorithm. It should be completely evident. We have seen that this immediately generalizes to any number of drawers and most general forms of quantum correlation.
The above is also confirmed by the fact that the time-symmetrization of the ordinary quantum description of the quantum algorithms says the same thing and is mandatory to satisfy the atemporal symmetry implicit in the latter description.
Summarizing, the teleological character of quantum algorithm should be evident, and enough to dispel the belief that teleological evolutions have no physical basis.
Appendix
We have seen that, according to the present theory, the usual quantum description of quantum algorithms is incomplete. How to complete it has already been discussed at a conceptual level in the previous part of the work. We now provide the formalization. We do it in the simplest instance of Grover algorithm; we shall see that everything immediately generalizes.
We should perform the following steps:
(1) Extend the usual quantum description of the quantum algorithm, limited to the process of solving the problem, to the upstream process of setting the problem: One should measure the problem-setting in a quantum superposition of all the possible problem-settings. The unitary computation and measurement of the corresponding solution follow.
(2) Relativize the extended quantum description of the quantum algorithm with respect to the problem-solver, from whom the problem-setting selected by the initial measurement should be concealed (it would tell her also the solution of the problem, before she begins her problem-solving action). Note that the advanced knowledge of one of the possible halves of the information about the solution of the problem must be in the quantum description of the quantum algorithm with respect to the problem-solver.
(3) Complete the quantum description of the quantum algorithm obtained so far by time-symmetrizing it.
Steps (1) through (3) are developed in the following correspondingly numbered sections.
(1) Extending the usual quantum description of quantum algorithms to the process of setting the problem
The usual quantum description of quantum algorithms is limited to the unitary process of computing the solution of the problem and the final measurement of the solution. We extend it to the upstream process of setting the problem by measuring the problem-setting in a quantum superposition of all the possible problem settings.
We introduce the notation; we make reference to the simplest, four drawer, instance of Grover algorithm:
We number the four drawers in binary notation: $00,01,10,11$.
We will need two quantum registers of two quantum bits each:
A register $B$ under the control of the problem-setter Bob contains the problem-setting observable $\hat{B}$ (the number of the drawer in which the ball is hidden) of eigenstates and corresponding eigenvalues:
$|00\rangle_{B}$ and $00$; $|01\rangle_{B}$ and $01$; $|10\rangle_{B}$ and $10$; $|11\rangle_{B}$ and $11$.
Measuring $\hat{B}$ in a quantum superposition (or, indifferently, mixture) of all the possible problem-settings selects one of them at random.
A register $A$ under the control of the problem-solver Alice contains the number of the drawer opened by her. The content of register $A$ is the observable $\hat{A}$ of eigenstates and corresponding eigenvalues:
$|00\rangle_{A}$ and $00$; $|01\rangle_{A}$ and $01$; $|10\rangle_{A}$ and $10$; $|11\rangle_{A}$ and $11$.
Measuring $\hat{A}$ at the end of the quantum algorithm yields the number of the drawer with the ball selected by the initial measurement, namely the solution of the problem. Note that $\hat{B}$ and $\hat{A}$ commute.
The quantum description of the extended quantum algorithm is:
$\begin{array}[]{ccc}\begin{array}[]{c}\;\text{time }t_{1}\text{, meas. of}\;\hat{B}\end{array}&t_{1}\rightarrow t_{2}&\text{time }t_{2}\text{, meas. of}\;\hat{A}\\ &&\\ \begin{array}[]{c}\left(|00\rangle_{B}+|01\rangle_{B}+|10\rangle_{B}+|11\rangle_{B}\right)\\ \left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)\end{array}&&\\ \Downarrow&&\\ |01\rangle_{B}\left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)&\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow&|01\rangle_{B}|01\rangle_{A}\end{array}$ (Table I)
Here and in the following we disregard normalization.
We should start from the quantum state in the upper-left corner of Table I diagram and then follow the vertical and horizontal arrows. Initially, at time $t_{1}$, both registers, each by itself, are in a quantum superposition of all their basis vectors. By measuring $\hat{B}$ (the content of $B$) in this initial state, Bob selects the number of the drawer with the ball at random; say it comes out $01$: see the state of register $B$ under the vertical arrow [^4] By the way, Bob would be free to unitarily change at will the sorted out number of the drawer with the ball. We omit this operation that would change nothing here. .
It is now Alice’s turn. The unitary part of her problem-solving action—by the unitary transformation $\hat{\mathbf{U}}_{1,2}$—produces the solution of the problem, of course, never changing the problem-setting.
Note that we will never need to know $\hat{\mathbf{U}}_{1,2}$ (i.e. Grover algorithm). We will only need to know that, between the outcome of the initial measurement and that of the final measurement, there is a unitary evolution that builds up one-to-one correlation between the two initially unrelated observables $\hat{B}$ and $\hat{A}$.
At the end of Alice’s problem-solving action, at time $t_{2}$, register $A$ contains the solution of the problem, i.e. the number of the drawer with the ball—bottom-right corner of the diagram.
Eventually, by measuring $\hat{A}$ (the content of register $A$), which is already in one of its eigenstates, Alice acquires the solution with probability 11—the quantum state in the bottom-right corner of the diagram naturally remains unaltered throughout this measurement.
(2) Relativizing the extended quantum description of the quantum algorithm with respect to the problem-solver
We must relativize the quantum description of the quantum algorithm of Table I, which is with respect to the customary external observer, with respect to the problem-solver. This amounts to concealing from her the projection of the quantum state associated with the initial measurement of the problem-setting
Since Alice’s problem-solving action never changes the number of the drawer with the ball selected by the initial measurement and the observables $\hat{B}$ and $\hat{A}$ commute, describing this concealment is simple. It suffices to postpone after the end of Alice’s problem-solving action the projection of the quantum state associated with the initial measurement of $\hat{B}$ (or, indifferently, to postpone the very measurement of $\hat{B}$). The result is the following quantum description of the quantum algorithm with respect to Alice:
$\begin{array}[]{ccc}\begin{array}[]{c}\text{time}\text{ }t_{1}\text{, meas. of}\;\hat{B}\end{array}&t_{1}\rightarrow t_{2}&\text{\text{time }\text{}}t_{2}\text{, meas. of}\;\hat{A}\\ &&\\ \begin{array}[]{c}\left(|00\rangle_{B}+|01\rangle_{B}+|10\rangle_{B}+|11\rangle_{B}\right)\\ \left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)\end{array}&\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow&\begin{array}[]{c}|00\rangle_{B}|00\rangle_{A}+|01\rangle_{B}|01\rangle_{A}+\\ |10\rangle_{B}|10\rangle_{A}+|11\rangle_{B}|11\rangle_{A}\end{array}\\ &&\Downarrow\\ &&|01\rangle_{B}|01\rangle_{A}\end{array}$ (Table II)
To Alice, the initial measurement of $\hat{B}$ leaves the initial state of register $B$ unaltered. At the beginning of her problem-solving action she is completely ignorant of the number of the drawer with the ball selected by Bob—see the state of register $B$ on the left of $\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow$.
The unitary transformation $\hat{\mathbf{U}}_{1,2}$ is then performed for a quantum superposition of the four possible numbers of the drawer with the ball. The state at the end of it—on the right of the horizontal arrows—is a quantum superposition of tensor products, each a possible number of the drawer with the ball in register $B$ tensor product the corresponding solution (that same number) in register $A$. Since $\hat{B}$ and $\hat{A}$ commute, the final measurement of $\hat{A}$ must project this superposition on the tensor product of the number of the drawer with the ball initially selected by Bob and the corresponding solution—bottom-right corner of Table II diagram.
Note that this relativization relies on relational quantum mechanics [26, 27] where the quantum state is relative to the observer. In the following, we will call this quantum description of the quantum algorithm with respect to the problem-solver its ordinary quantum description.
(3) Time-symmetrizing the ordinary quantum description of the quantum algorithm
As anticipated in the Introduction, the ordinary quantum description of Grover algorithm is incomplete and is completed by time-symmetrizing it. We should require that the selection of the information that specifies the sorted out pair of correlated measurement outcomes, the problem-setting and the corresponding solution, evenly shares between the initial measurement of the problem setting and the final measurement of the solution. The half information selected by the initial measurement should propagate forward in time by the unitary computation of the solution $\hat{\mathbf{U}}_{1,2}$ until becoming the state immediately before the final measurement. Time-symmetrically, the complementary selection performed by the final measurement should propagate backward in time by the inverse of the computation of the solution $\hat{\mathbf{U}}_{1,2}^{{\dagger}}$ until replacing the previous outcome of the initial measurement.
We consider the time-symmetrization instance where the initial measurement of $\hat{B}$ reduces to the measurement of $\hat{B}_{l}$, the left bit of the two-bit number contained in register $B$; say that it randomly selects left bit 0. The final measurement of $\hat{A}$ should correspondingly reduce to that of $\hat{A_{r}}$, the right bit of the two-bit number contained in register $A$; say that it randomly selects right bit $1$. The randomly selected number of the drawer with the ball is thus $01$.
In this particular time-symmetrization instance, the ordinary quantum description of the quantum algorithm of Table II becomes:
$\begin{array}[]{ccc}\begin{array}[]{c}\;\text{time }t_{1}\text{, meas. of}\;\hat{B}_{l}\end{array}&t_{1\rightleftarrows}t_{2}&\text{time }t_{2}\text{,meas. of}\;\hat{A_{r}}\\ &&\\ \begin{array}[]{c}\left(|00\rangle_{B}+|01\rangle_{B}+|10\rangle_{B}+|11\rangle_{B}\right)\\ \left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)\end{array}&\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow&\begin{array}[]{c}|00\rangle_{B}|00\rangle_{A}+|01\rangle_{B}|01\rangle_{A}+\\ |10\rangle_{B}|10\rangle_{A}+|11\rangle_{B}|11\rangle_{A}\end{array}\\ &&\Downarrow\\ \begin{array}[]{c}\left(|01\rangle_{B}+|11\rangle_{B}\right)\\ \left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)\end{array}&\Leftarrow\hat{\mathbf{U}}_{1,2}^{{\dagger}}\Leftarrow&|01\rangle_{B}|01\rangle_{A}+|11\rangle_{B}|11\rangle_{A}\end{array}$ (Table III)
The projection of the initial state of register $B$ due to the measurement of $\hat{B}_{l}$—top-left corner of the diagram—is postponed after Alice’s problem-solving action, namely outside the present table which is limited to that action. With this postponement, the initial state of register $B$ goes unaltered through the initial measurement of $\hat{B}_{l}$ becoming, by $\hat{\mathbf{U}}_{1,2}$, the quantum superposition of tensor products above the vertical arrow. Alice’s final measurement of $\hat{A_{r}}$, which in this time-symmetrization instance selects the value $1$ of the right bit of register $A$, projects the state above the vertical arrow on the state below it. Then (in causal order), this final measurement outcome propagates backward in time by $\hat{\mathbf{U}}_{1,2}^{{\dagger}}$ until becoming the definitive outcome of the initial measurement. This backward in time propagation is an instance of the time-symmetrized quantum algorithm. As we can see from the diagram, the new outcome of the initial measurement tells the problem-solver, before she begins her problem-solving action, one of the possible halves of the information about the solution, namely that the ball is in drawer either 01 or 11: look at the state of register $B$ in the bottom left corner of the table.
Note that, for the mathematical equivalence between $\Leftarrow\hat{\mathbf{U}}_{1,2}^{{\dagger}}\Leftarrow$ and $\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow$, this bottom line of Table III diagram (an instance of the time-symmetrized quantum algorithm) mathematically can also be read in the usual left to right way:
$\begin{array}[]{ccc}\text{time }t_{1}&t_{1}\rightarrow t_{2}&\text{time }t_{2}\\ &&\\ \begin{array}[]{c}(|01\rangle_{B}+|11\rangle_{B})\\ \left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)\end{array}&\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow&|01\rangle_{B}|01\rangle_{A}+|11\rangle_{B}|11\rangle_{A}\end{array}$ (Table IV)
We can see that, in each instance of the time-symmetrized quantum algorithm, the computational complexity of the problem to be solved by Alice quadratically reduces from the problem of locating the ball in one of $4$ drawers to that of locating it in $2=\sqrt{4}$ drawers. In still equivalent terms, the dimension of the Hilbert space in which the search for the solution takes place quadratically reduces from $4$ basis vectors to $2=\sqrt{4}$ basis vectors.
In summary, the causal loop generated by the time-symmetrization process is the zigzag diagram of Table III. We can see that it is one of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup. The state of register $B$ in the top-left part of the diagram tells us that Alice, before beginning her problem-solving action, is completely ignorant of the number of the drawer with the ball. Through the zig-zag, this state changes into the state of register $B$ in the bottom-left part of the diagram, which tells us that Alice, before beginning her problem-solving action, knows in advance that the ball is in drawer either $01$ or $11$; we are dealing with an effective causal loop that changes its own past. She can then solve the problem in an optimal logical way by opening just one of the two drawers, indeed as in the classical logic description of what the quantum algorithms does.
Also note that table III diagram goes along with the fact that the unitary part of the quantum algorithm, $\hat{\mathbf{U}}_{1,2}$, takes the opening of just one drawer, but does not imply it. We know this fact from outside Table III diagram: we know that this is what the unitary part of Grover algorithm takes. The diagram says that the problem-solver knows in advance that the ball is in a pair of drawers rather than four—bottom left corner of the diagram—and this could allow her to solve the problem by opening a single drawer—bottom line of the diagram of table III or the diagram of table IV. The advanced knowledge rule tells us the room for the quantum speedup but does not imply it.
Eventually, we should take the quantum superposition of all the time-symmetrization instances.
Note that this time-symmetrization, mathematically, leaves the unitary computation of the solution, $\hat{\mathbf{U}}_{1,2}$, unaltered, but changes the behavior of causality along it. The single causal loop of Table II is replaced by the quantum superposition of the causal loops whose mutually exclusive or is logically implied by the sheer existence of the quantum speedup.
In subsection 2.6, we have seen that all this, first, immediately extends to any number of drawers, then, to more general forms of quantum correlation.
Acknowledgements
Thanks are due to Eliahu Cohen, Artur Ekert, Avshalom Elitzur and David Finkelstein for useful comments, Daniel Sheehan for organizing the San Diego AAAS Pacific Division conferences on retrocausality, a far-looking forum for the discussion of frontier, also unorthodox, physics, and Mario Rasetti for organizing the first series of international workshops on quantum communication and computation, the 1991–96 Elsag Bailey–ISI Turin Villa Gualino workshops.